### v3

# Jules Drean MIT CSAIL Miguel Gomez-Garcia MIT CSAIL Fisher Jepsen MIT CSAIL Thomas Bourgeat EPFL  and  Srinivas Devadas MIT CSAIL

Abstract.
Enclaves or Trusted Execution Environments are trusted-hardware primitives that make it possible to isolate and protect a sensitive program from an untrusted operating system. Unfortunately, almost all existing enclave platforms are vulnerable to microarchitectural side channels and transient execution attacks, and the one academic proposal that is not does not allow programs to interact with the outside world. We present Citadel, to our knowledge, the first enclave platform with microarchitectural isolation to run realistic secure programs on a speculative out-of-order multicore processor. We show how to leverage hardware/software co-design to enable shared memory between an enclave and an untrusted operating system while preventing speculative transmitters between the enclave and a potential adversary. We then evaluate our secure baseline and present further mechanisms to achieve reasonable performance for out-of-the-box programs. Our multicore processor runs on an FPGA and boots untrusted Linux from which users can securely launch and interact with enclaves. To demonstrate our platform capabilities, we run a private inference enclave that embed a small neural network trained on MNIST. A remote user can remotely attest the enclave integrity, perform key exchange and send encrypted input for secure evaluation. We open-source our end-to-end hardware and software infrastructure, hoping to spark more research and bridge the gap between conceptual proposals and FPGA prototypes.
1. Introduction
Secure Enclaves or Trusted Execution Environments (TEEs) (costan2016sanctum, ; aegis, ; ferraiuolo2017komodo, ; lee2020keystone, ; bahmani2021cure, ; feng2021scalable, ), are trusted-hardware primitives used to protect the execution of sensitive programs. They make it possible to run sensitive computations in isolation with strengthen privacy and integrity properties. Academic TEE proposals date back twenty years to XOM (xom, ) and AEGIS (aegis, ) and industry implementations include ARM TrustZone (trustzone04, ; pinto2019demystifying, ), Intel SGX (mckeen2013sgx, ; costan2016intel, ) and AMD SEV (sev2020strengthening, ).
Unfortunately, these kinds of platforms have been shown to be vulnerable to microarchitectural side channels and transient execution attacks (chen2019sgxpectre, ; koruyeh2018spectre, ; ryan2019hardware, ; van2020lvi, ). In particular, attacks such as those in (van2018foreshadow, ; van2020sgaxe, ) were able to not only extract sensitive data from Intel SGX enclaves, but also managed to break the security of the remote attestation mechanism.
One academic proposal (bourgeat2019mi6, ) has looked at securing enclaves in the context of microarchitectural side channels and transient execution attacks. They aim for a strong security guarantee where a potential attacker can only observe the enclave’s output. Unfortunately, this extreme isolation strongly restricts the class of programs that can be run, as no communication with the outside world is allowed except when entering or exiting the enclave.
Figure 1. Citadel’s Computing Stack
We are not the first to observe that restrictive enclave execution models limit the programmability and the performance of enclave programs(van2023cheri, ; yu2022elasticlave, ; li2021confidential, ; lee2022cerberus, ; park2020nested, ; steinegger2021servas, ) but those work ignore security challenges introduced by microarchitectural side-channels and transient execution attacks. When considering those challenges, even basic communication primitives like public shared memory with the operating system are not straightforward to implement efficiently and securely.
In this paper, we try to fill the gap between platform that have strong security but restrict enclave programs to the extreme and platforms that allow better enclave execution models but ignore side channels and transient execution attacks.
We first define our relaxed security goal: an attacker should only be able to observe the transcript of non-speculative accesses to a shared memory region. We then craft requirements for the hardware and the software so that, together, they can prevent secrets from leaking through speculative access to shared memory. In practice, the hardware will isolate private microarchitectural state and restrict the victim program speculation model to help it achieve the just described security guarantee. Intuitively, the more restrictive the speculation model, the lighter the requirements on the program, but the least performance we get.
In our case, we first consider a very conservative speculation model where no speculative shared accesses are allowed. This effectively prevents the existence of any speculative transmitters between the victim and the attacker. We describe how to enable shared memory securely for the specific open-source out-of-order design we used, while disabling speculative shared accesses. We also make sure to not enable any new speculative transmitters by bypassing the L1 D cache, and to not break address translation by tagging the TLB’s and the Address Translation Cache’s entries. Finally, we give a brief overview of the software infrastructure we have implemented in order to boot Linux and run enclaves from user programs on our FPGA prototype.
We use this secure baseline to evaluate the performance of the kind of programs that one could be interested in putting in an enclave. Out-of-the-box, we obtain an underwhelming $120.6\%$ performance overhead for a cryptographic library. We identify and analyze the performance bottlenecks to design a series of hardware-software mechanisms to improve both enclaves’ and unprotected processes’ performance. More precisely, we develop a reconfigurable Last Level Cache (LLC) partitioning scheme, a performance necessity for our small LLC. This reduces our overhead from $120.6\%$ to $19.4\%$. We also introduce software controls to enable a slightly less conservative enclave speculation model, where pipelined accesses to shared memory are allowed but predictors are hardwired to $pc+4$. The efficient way to use this mode is on code segments with burst accesses to shared memory. Under this less conservative speculation model, the programmer needs to make sure that straight-line speculation ($pc+4$) within these segments will not leak any secrets. This mechanism reduces overheads from $19.4\%$ to $2.7\%$ on our cryptographic benchmark.
Our final prototype is a modified out-of-order RISC-V 2-core processor synthesized to run on an AWS F1 FPGA. It fully implements all our security and dynamic resizing mechanisms, boots (untrusted) Linux and makes it possible for user-level applications to launch and safely communicate with secure enclaves executing real applications such as a lightweight Python runtime, a cryptographic library or a private inference application running a small neural network. We also implement an end-to-end attestation mechanism, making it possible for an enclave to identify itself independently of where the OS placed it in physical memory, enabling enclaves to act as a trusted third party. We implement and make all the pieces of software and modified hardware available, such as our secure bootloader, our security monitor and a Linux kernel module to interact with the security monitor. We summarize the different elements implemented for this paper in Figure 1. This makes our platform a unique environment for research in the domain of microarchitectural side channels, transient execution attacks and secure enclaves. We hope researchers will take ownership of the platform, attack it, modify it to prototype and test complex hardware-software mechanisms and research ideas.
To summarize, the main contributions of the paper are:
  • To our knowledge, we show the first end-to-end enclave platform running securely on an FPGA implementation of a speculative out-of-order processor.
  • We show how to build a shared memory mechanism while guaranteeing an attacker will only observe non-speculative accesses to shared memory.
  • Guided by early performance evaluation, we propose a series of hardware-software mechanisms to make the platform more configurable and improve both enclaves’ and unprotected processes’ performance.
We open-source our hardware, software and debugging infrastructure at https://github.com/mit-enclaves/citadel.
2. Motivation & Background
We want to : 1) run programs that can interact with the outside world. 2) formalize the guarantees that we expect the hardware to provide 3) have good performance.
2.1. Security Monitor
In many enclave platforms, including Citadel, the Security Monitor (SM) (costan2016sanctum, ) (lebedev2019sanctorum, ) is a small (9K LOC), trusted piece of software running at a higher privilege mode than the hypervisor or the OS. Its role is to link low-level invariants exposed by the hardware (e.g., if a specific memory access is authorized), invoking the right microarchitectural cleanup routine, and the high-level security policies defined by the platform (e.g., which enclave is currently executing on which core and which memory region does it own).
2.2. Microarchitectural Timing Side Channels
Microarchitectural side channels allow for information to be leaked from one security domain to another. These attacks exploit microarchitectural structures shared between different security domains, such as memory caches (gotzfried2017cache, ; brasser2017software, ), translation look-aside buffers (TLBs) (gras2018translation, ), branch predictors (evtyushkin2018branchscope, ; aciiccmez2006predicting, ), DRAM controllers (wang2014timing, ), etc., to transmit a signal. Because microarchitectural state cannot be directly observed through the Instruction Set Architecture (ISA), an attacker will typically rely on the timing of events to indirectly observe the microarchitectural state. For instance, if a piece of data has been loaded in the cache, accessing it will be faster than if it is only stored in main memory. Using various methods to prepare the shared microarchitectural state and observe the state changes due to the victim’s activity, the attacker can recover sensitive information about the victim’s execution, for example, which addresses have been accessed, which branches were taken in which directions, and/or information about victim program’s inputs. Cache side channels are the easiest channels to exploit, and so have been extensively studied, but any piece of shared microarchitectural state can theoretically be used as well.
To summarize the observations that could be made through side-channels, the community formalizes side-channels as a set of observations that an attacker could make about the microarchitectural state. For example, cache side-channels on unprotected machines typically lead to an observation model that includes the sequence of addresses of memory operations (sometimes with their timestamps) as part of the observation.
2.3. Transient Execution Attacks and Execution Models
Speculative or transient execution is one of the key performance features of modern processors. Unfortunately, it can make reasoning about security boundaries more difficult. Some microarchitectural data structures that orchestrate speculation (e.g., branch predictors) can be trained so the processor will execute code snippets speculatively, in ways that are incorrect or on data that should not be accessed during normal execution. Concretely, instructions that were not part of the program normal execution path can be executed out-of-thin-air, speculatively, by the processor. An example is the execution of code speculatively before the resolution of a preceding security check, such as a memory access bound. These erroneous executions happen speculatively, and these instructions will be squashed by the processor. Nonetheless, they will still change the microarchitectural state of the machine. For example, a speculative load to some data will not affect the values of the architectural registers but will load the data in the memory hierarchy and possibly change the state of caches. This means that a program that manages sensitive data can be coerced into leaking private data speculatively through a side channel. More importantly, once some specific code snippet, or gadget, has been identified, these attacks can be used to leak virtually any data located in the victim’s virtual address space. This type of attack, called Spectre, was first introduced in (Kocher2018spectre, ), and many variants have since been demonstrated.
3. Defining Security
To achieve security, we rely on both the hardware to enforce some isolation and on the fact the software verifies some requirements: we do not protect all-software out-of-the-box.
Intuitively, for a given security property, if the isolation enforced by the hardware is stronger, then the requirements on the software are weaker and we can enforce the security property for more programs.
Previous Work Previous work looking at enclaves in a context of transient execution attacks (bourgeat2019mi6, ) aims at limiting the attacker’s view to the enclave final output. Completion time was considered outside of the threat model and could be independently handled using traffic shaping techniques (padding) techniques. Unfortunately, this security property is too restrictive as it only considers a limited class of programs that do not communicate with the outside world.
Public Shared Memory We define a weaker security property in order to consider a more important class of programs to run securely. Our goal is for the system to guarantee that an attacker should only be able to observe the transcript of non-speculative accesses to a shared memory region. In particular, we don’t want an attacker to be able to observe any speculative accesses to shared memory from the victim nor any other microarchitectural state (accesses to private memory, branch predictors, TLBs etc…). However, similarly to previous work, we do not consider the timing of each access is observable by the attacker, i.e., there are no timestamps for the non-speculative shared memory accesses. We consider that leakage through the timing of shared memory accesses should be addressed using orthogonal traffic-shaping techniques such as (deutsch2022dagguise, ; wang2014timing, ; shafiee2015avoiding, ; zhou2017camouflage, ) and is left for future work.
Hardware & Software Requirements We now define what kind of requirements the hardware will needs to enforce and how the hardware restricts the victim program speculation model to help it achieve the just described security guarantee. 1) the hardware should properly isolate private microarchitectural state. 2) together, the hardware and the software need to prevent secrets from leaking through speculative accesses to shared memory. Intuitively, the more restrictive the speculation model of shared memory accesses enforced by the hardware, the lighter the conditions on the program to finally achieve security, but the least performance we get.
We first experiment with a restrictive speculative model where the victim program is not able to perform speculative accesses to shared memory. This effectively prevents the existence of any speculative transmitters between the victim and the attacker. The program only needs to make sure that the explicit shared memory accesses don’t leak information about the program secrets. Any constant-time cryptographic code satisfies these restrictions. Unfortunately this restrictive model gives us poor performance.
We experiment with a second speculative model called Burst mode where speculation on shared memory accesses is allowed, but all the predictors are deactivated and hard-coded to $pc+4$. The rationale is to allow pipelined accesses to shared memory to improve performance. Burst mode is intended to be only activated for code segments with straight-lined, burst accesses to shared memory. Under this less conservative speculation model, the software requirements are more important and the programs needs to verify that straight-line speculation ($pc+4$) within these segments will not leak any secrets. Outside of the said segments, we use the first speculation model.
3.1. Attacker Capabilities
We now complete our threat model by describing the attacker’s capabilities. The attacker is collocated on the same machine as the victim and can mount any software-based attack, including microarchitectural side channels and transient execution attacks. The attacker has compromised the majority of the software stack, including the operating system (OS) except for a limited trusted code base (TCB) including the secure bootloader, the SM (together 15K lines of code (LOC)) and the enclave. For the rest of this paper, we refer to any code outside of the TCB as the OS. We ignore attacks requiring physical access to the machine such as physical (e.g., noise, electromagnetic field, etc.) side channels. We also ignore side channels that can be mounted in software, but exploit the physical layer like power side channels. We do not consider Rowhammer attacks. As highlighted by our security definition, we are worried about enclave privacy and do not consider denial of service.
4. Citadel 1.0: A Secure Baseline
Even though defense mechanisms often propose to ”not speculate” when executing some instructions that can dirty the microarchitectural state, implementing such policies is not completely straightforward. In this section, we describe how to securely enable shared memory for our specific out-of-order open-source design while disabling speculative shared accesses. We also make sure to not enable any new speculative transmitters by bypassing the L1 D cache, and to not break address translation by tagging the TLB’s and the Address Translation Cache’s entries. Finally, we give a brief overview of the software infrastructure we have implemented in order to boot Linux and run enclaves from user programs.
4.1. Hardware Isolation
Memory Isolation Sanctum (costan2016sanctum, ) already presents hardware mechanisms for memory isolation in the context of an untrusted OS. Like RISC-V Physical Memory Protection (PMP) (waterman2015risc, ), Sanctum introduces a mechanism to partition physical memory and prevent a malicious OS from accessing another security domain’s private memory regions, or different enclaves from accessing each other’s private memory. At its core, the mechanism uses two 64-bit registers per core (one for trusted memory and one for untrusted/shared memory) as memory bitmaps. Each bit represents a fixed-size physical memory region (typically 32MB) and functions as an access-rights flag for the currently-running program. Citadel borrows this mechanism for its simplicity and enforces these checks on every memory access, including speculative accesses and for accesses performed during page walks.
Microarchitectural Isolation Following MI6 (bourgeat2019mi6, ) methodology we partition all spatially shared microarchitectural state (LLC, MSHRs…) and flush temporally shared state on context switch (all cores’ pipeline state, L1 caches, TLBs…). We are not vulnerable the DRAM controller side channels as we inherit MI6’s constant time DRAM padding mechanism. Finally, our processor does not support hyper-threading and we are not vulnerable to related side channels such as port contention.
4.2. Enabling Shared Memory
Figure 2. The Dual Page Table Mechanism
Sanctum also uses a dual-page-table mechanism (see Figure 2) to differentiate between trusted and untrusted memory based on virtual addresses, which we also borrow in Citadel. Two new registers (equivalent to base and bounds) are used to define the trusted virtual memory range (i.e. the enclave’s virtual memory). The SM enforces the invariant that the physical memory mapped as trusted is owned by the corresponding enclave. When a memory access to a virtual address is performed by a program, one of the two page tables needs to be picked to translate the virtual address. If the virtual address falls within the trusted range, the translation is performed using the enclave page table. Another important invariant enforced by the SM is that the enclave page table is always stored within enclave private memory. If the virtual address falls outside the trusted range, it corresponds to an untrusted or shared memory access and the translation is performed using the OS page table. The OS page table is untrusted and stored in OS memory. Note that for every physical address used during address translation, additional bound checks are performed to ensure they belong to the same security domain (enclave or OS memory) as the translated address. This mechanism is especially handy to eliminate speculative transmitters, as there is no need to perform any speculative address translation (which could leak information) to know whether a memory access is trusted or not.
4.3. Restricting Speculative Access to Shared Memory
In the baseline out-of-order processor we are hardening, memory instructions are decoded, renamed and then simultaneously queued in the reorder buffer (ROB) and the memory reservation station (MemRS). When an instruction’s operands are ready, it is dispatched to the memory execution pipeline (MemExePipeline). Our approach identifies instructions that access untrusted memory early in the MemExePipeline and delay the translation and processing of the memory accesses until the instruction is no longer speculative, i.e., when it reaches the head of the ROB. These changes require the modification of MemRS and MemExePipeline.
Figure 3. Changes to Memory Execution Pipeline. New elements are shown in blue.
Modifying the Memory Execution Pipeline Upon entry into the MemExePipeline, checks are performed on memory instructions to decide if the memory operation is safe to translate and then issue, or not. It is important to perform these checks very early as the next steps in the MemExePipeline touch shared microarchitecture (through address translation and cache hierarchy for instance) potentially transmitting information to an attacker. A memory access is deemed safe if it accesses trusted memory OR if it is non-speculative. Because the enclave trusted memory range is defined in virtual memory, this check can be performed as soon as operands (i.e. virtual address) are available. If an instruction tries to access untrusted or shared memory, we check that it is non-speculative by conservatively checking if it has reached the head of the ROB. If the instruction is safe, it is dispatched to the next stage of the MemExePipeline. If it is not safe, the instruction is squashed and a signal is sent to the MemRS indicating that it should re-dispatch it once it reaches the head of the ROB.
Modifying the Memory Reservation Station Our mechanism requires MemRS being able to re-dispatch some instructions once they become non-speculative: untrusted memory loads will indeed typically need to get dispatched twice.
Disabling the Mechanism for the OS Blocking speculation on memory access can have a significant performance impact. The SM makes sure that nonenclave programs (typically the OS and standard usermode programs) are still allowed to speculate for their memory accesses by disabling the trusted memory mechanism.
4.4. Speculative Transmitters in the L1 D Cache
The core-private part of the memory system, including the L1 D cache, is now shared between trusted memory accesses and untrusted or shared memory accesses. While shared memory cannot be accessed speculatively, trusted memory can. It means that shared memory can be speculatively evicted by a speculative trusted memory access. Furthermore, this eviction will be observable to an attacker. The cache coherency protocol is such that writing to a memory location that is cached by another core will take more time than writing to a memory location that is not cached by another core. As a result, a shared L1 D can be leveraged by an attacker to create a speculative transmitter and mount a spectre-style attack. To our knowledge, we are the first to identify this subtle attack vector. Note that this doesn’t only apply to shared cache sets, but it also applies to contention over other shared resources in the L1 D cache (including the pipeline to access the SRAM). Partitioning the L1 D cache would be too costly, instead we bypass the L1 D cache for untrusted or shared accesses and directly send the requests to the LLC.
4.5. Tagged TLB for Correct Address Translations
When using the dual page table mechanism, the core-private part of the memory system is shared between trusted memory accesses and untrusted memory accesses. For functional correctness, we need to make sure that addresses are translated correctly for each domain. This small microarchitectural structure accelerates address translation by caching intermediate results used by the page walker. Unfortunately, only a few upper bits of the virtual address are used at the beginning of a page walk, easily resulting in aliasing across the two domains. We tag the Address Translation Cache by adding a bit to each entry indicating if the cached information applies to the trusted or untrusted memory address translations. Aliasing can also occur in the TLB if the trusted virtual memory range is smaller than the granularity of a virtual memory page (which can be if we use Mega or Giga Pages). We also tag the TLB entries to avoid address mistranslations.
4.6. Miscellaneous
We statically partition MSHRs between two cores and place a fair round-robin arbiter at the entry of the LLC. Because these two mechanisms only marginally affect the latency of the LLC for a 2-core design, we observe negligible ($\approx$ 0.1%) performance impact.
We also expose controls for speculation and predictors to the software through Control Status Registers (CSRs). Specifically, a program is able to deactivate speculation all together (every instruction waits to reach the head of the ROB), delay the issuing of shared memory accesses to the head of the ROB, deactivate the training of predictors (branch predictors and return address stack) and/or their use (and then it falls back to the predictor $pc+4$).
4.7. Building the Software Infrastructure
We are interested in evaluating our platform on a Linux-ready system and for programs that one would be interested in putting in an enclave (e.g. cryptographic libraries). This means a lot of software needs to be developed to be able to launch enclaves from a Linux user mode process.
First we implement the SM at the most privilege level of RISC-V using a mix of assembler and C and add the necessary pieces of a micro-kernel required to boot Linux on top of it. We also implement what we call a Mini-SM which is a downsized copy of the SM private to each enclave. This avoids contention over most SM resources between enclaves and untrusted programs. To protect the SM against transient execution attacks, we disable speculation during SM execution. Second, we develop a Linux Kernel module that allows user programs to request SM calls and set up enclaves, keeping Linux responsible for the resource management. Finally, we implement a Secure Bootloader and an end-to-end attestation mechanism that makes it possible to uniquely identify an enclave, establish a shared secret with it and exchange encrypted private data. Our attestation mechanism makes it possible for an enclave to be placed anywhere in physical memory and still have consistent identity through the measurement of the direct mapping between virtual addresses and memory content without measuring physical addresses. The unique mapping is attested and enforced by the SM during enclave setup. More implementation details can be found in the evaluation section (see section 6).
5. Citadel 2.0: Improving Performance
Our first baseline successfully achieves security. In this section, we show several mechanisms to improve its performance. First. fixed LLC partitioning á la MI6 destroys performance for our 2-core machine. We show a reconfigurable LLC partitioning scheme and a fine grain flushing mechanism. Furthermore, disabling speculative shared memory accesses makes software requirements especially light, but provides limited performances. We show a relaxed speculative model where we enable pipelined accesses to shared memory but disable predictors and hard-code them to $pc+4$. When this mode is enabled, the program needs to verify more requirements, i.e., that no secret is leaked when speculating $pc+4$. We show that this new mode improves performance for code segments with burst shared memory accesses, but it hampers performances in the general case. We observe that this is due to the TLB being trashed by misspeculations across functions context. This is an effect that was studied (in a nonsecurity context) in the early 2000s(onur1, ; onur2, ; onur3, ): misspeculations can induce large overhead not only due to the delay in executing the right path, large delays are induced due to the trashing of the TLBs. Finally, we show how insights from building the SM make it possible to securely optimize context switch and enable frequent interrupts.
5.1. Evaluation-Guided Optimisations
Our secure baseline runs on FPGA with a Linux environment (see section 6 for our full configuration) which makes it possible to analyses performance bottlenecks on real programs. For this section, we use an elliptic curves (Ed25519) cryptographic library as a running example (see section 6 for more details). At a high level, our enclave interactively receives messages to sign from a benchtest program using shared memory. Out-of-the box, we see an underwhelming overhead of 120% overhead compared to directly (and insecurely) linking the cryptographic library to our benchtest program.
5.2. Reconfigurable LLC Partitioning
An obvious bottleneck in our design is cache partitioning. Sanctum and MI6’s mechanism (i.e. dividing the LLC in 64 equal size slices to match the 64 DRAM regions) is a terrible idea for our system. Our 2-core prototype has a 1MB LLC (see 5), which divided by 64 gives us 16KB LLC slices. This is smaller than the 64KB L1 D+I caches. Because our caches are inclusive, this means an enclave or a program that resides in one memory region cannot even fully use the L1 caches. MI6 circumvented that issue by assuming a 16-core design with a large 16MB LLC.
We build a mechanism for the OS to reconfigure the LLC partitions to improve enclave and other program’s performance. This costly operation is only required when changing the LLC layout, which only happens before launching an application or an enclave with an important memory footprint. The OS requests a reconfiguration through an SM call while no enclaves are running. The SM will first check that the new partitioning is valid, doesn’t affect regions assigned to the SM. It then waits for other cores to flush their L1 caches and to enter a blocked state in the SM, making sure no requests will be sent to the modified regions during reconfiguration (which could put the cache in an incoherent state). The SM will then interact with the hardware to flush the LLC slices that are modified and set the LLC in its new configuration.
At the hardware level, we now design a practical reconfigurable LLC partitioning scheme. We cannot use way partitioning (kiriansky2018dawg, ) as the LLC has only 16 ways, and we need up to 64 slices, and Ironhide’s (ironhide, ) strategy assumes limited reconfiguration. We propose a new, simple, configurable set partitioning scheme that only requires minimal changes in the hardware.
The configuration for each LLC slice (base and bound of the set-index range) is stored in a big register (1216 bits) called the set-index range table. The SM can modify the entries of the set-index range table by sending MMIO requests to the cache, which is seen as a new device. When the LLC receives a request, it will compute the new set index for the corresponding address. We represent the transformation used in Figure 4. First, we extract a region ID from the address upper bits and obtain the corresponding index range base and size by looking up the set-index range table 1. We then extract the address original set index, located in the middle of the address, and tweak it. We compute the original set-index value modulo the size of the index range 2 and add the base index 3. This effectively restricts the new set index to the corresponding index range, partitioning the LLC 4. Note that we must extend the length of the tags used by the LLC to include the entire original index in case an index range is of size 1 and all addresses from the same region map to the same set.
We observe on our cryptographic library that we reduce overheads from 120% to 19.4% by using a non-uniform partition for the LLC. More details can be found in Section 6.5.
Figure 4. LLC set indexing scheme. The function takes an address as an input and output a set index in the LLC slice allocated to its memory region.
LLC fine-grain flush To go hand-in-hand with our reconfigurable LLC partitioning, we also want to be able to selectively flush parts of the LLC. We create a Zero Device in the upper physical address space. It is implemented as a simple piece of logic connected to the cache hierarchy at the same level as the DRAM, but that always sends back zeros. The trick is that the address space covered by the zero device has the same size as the DRAM and is located at addresses that will collide with DRAM addresses in the cache. Once the SM has placed the machine in a correct waiting state (where all the cores have flushed their L1 and are waiting in the SM), it only needs to access an eviction set in the zero-device address space to flush the cache slice. This means that flushing a small LLC slice is much faster than flushing a big one. We look at the performance of our flushing mechanism in the evaluation section (see 6.5).
5.3. Pipelined Shared Accesses
At this step, we observe that our test benchmark still present a 19.4% overhead compared to our insecure benchmark. We decide to measure the performance of different sections of the code to find and analyze the bottleneck. We identify the function creating the slowdown to be located inside the SHA256 library. This function performs burst accesses to shared memory to load the message to be signed. This behavior is exacerbated by the compiler, trying to optimize the code by placing all the load instructions to shared memory one after the other. Indeed, with our current mechanism, memory accesses to shared regions, first cannot use the L1 and have a 10 cycle latency, second cannot be pipelined as each instruction needs to wait to reach the head of the ROB to be sent. This is especially costly during burst accesses. The LLC supporting 16 outstanding requests and having a 10 cycles latency, 16 memory accesses should be pipelined to take a total of $10+16=26$ cycles. Without pipelining, they will now take $10*16=160$ cycles.
As a result, we designed a mode where we temporarily change the speculation model to allow shared memory burst accesses (as pipelining is a form of speculation). Nevertheless, we don’t want to do so while enabling any type of speculation. Indeed, we don’t want our program to speculatively jump to a spectre-like gadget and transmit secrets to our attacker. To avoid this, we tweak the speculation model to allow pipelined accesses to shared memory but disable the use and training of any predictors that could result in the speculative execution of rogue code. Specifically, we disable the Branch Target Buffer (BTB), the Branch History Table (BHT) and the Return Address Stack (RAS). Effectively, that guarantees that the hardware will always predict that the next program counter to be “$pc+4$” and avoid speculative execution of arbitrary code. We call this straight-line speculation.
Modifying the speculative model also modifies the hardware software contract. The programmer now needs to ensure that when this mode is enabled, no speculative transmitter can be reached by the program when doing straight-line speculation. Static analysis can be done manually or by reusing many existing tools: e.g. oo7(wang2019oo7, ), Kasper(johannesmeyer2022kasper, ), Kleespectre(wang2020kleespectre, ), Spectector(guarnieri2020spectector, ). Note that the usual critique and limitation of these tools are that 1) the code bases are too big to analyze without producing a lot of false positives 2) they make simplifying assumptions on the speculation model. Interestingly, both of those limitations are not concerns in the context of Citadel’s burst mode because the code snippets we analyze are small (e.g. boiling down to a memcpy function for instance) and the speculation model is well-defined, simple and enforced (only speculates PC+4). Hence, these works are well-suited to our needs. We consider adapting these tools to our more simple setup as future work. We add two new configurations in our software-exposed configuration registers to enable and disable this “burst” mode. The right way to use this mode is, like in our case, during burst accesses to shared memory (e.g. a memcpy function moving data from public memory to private enclave memory). The programmer only needs to identify these series of burst access and guard them with instructions activating and deactivating burst mode. When doing so on our cryptographic benchmark, we observe that overhead falls from $19.4\%$ to $2.7\%$. More results are presented in the evaluation section.
5.4. Relaxed Flushing Policies
Because we build the entire software infrastructure required to run enclaves, we were able to observe that some security domains, especially the enclave and some parts of the mini-SM, do not have secrets to keep from each other. Specifically, the mini-SM never manipulates sensitive data. As a result, the mini-SM does not need to flush microarchitectural state when context switching to enclaves. This reduces the cost of mini-SM call and opens the door to the use of frequent software or timer interrupts. To disable flushing selectively, we modify the hardware to deactivate systematic flushing and exposed to the SM a register to trigger the flush. We write the SM to trigger the flushing mechanism on required context switches, but not on most traps and SM-calls from the enclave. We evaluate the performance implication of our mechanism in Section 6.6.
6. Evaluation
6.1. Experimental Setup
Frontend
2-way superscalar, 256 entries BTB, 8 entries RAS, tournament predictor
Reorder Buffer
64 entries, 2-way insert/commit
Reservation Stations
2 stations for ALU (16 entries), 1 for FPU (16 entry), 1 for memory (16 entries)
L1 (I/D)
32KB, 8 ways, 8 outstanding requests
Load-Store Unit
24 LdQ entries, 14 StQ entries, 4 SB entries
LLC (L2)
1MB, 10 cycles latency, 16 ways, max 16 outstanding requests
Memory
120 cycles latency (24 outstanding requests)
Figure 5. Citadel’s microarchitectural configuration.
Citadel’s processor is a modified out-of-order 2-core based on Riscy-OOO (zhang2018composable, ) and MI6 (bourgeat2019mi6, ). The parameters used for the cores and the memory subsystem are given in Figure 5. Citadel runs on an AWS EC2 F1 instance, equipped with a Virtex UltraScale+ FPGA. On this FPGA board, Citadel can be clocked at $30$ MHz, which is similar to the baseline design. We use the RISC-V tool chain with GCC 12.2.0. We compile all software with the -O3 flag. We run Linux Kernel 6.2.0. Unless mentioned differently, all performance numbers are collected by running the benchmark on the FPGA using performance counters and an average of ten runs.
6.2. Implementation Data and TCB Breakdown
Component
LOC Size
Processor (bourgeat2019mi6, ) (+ modification)
$\sim 60$ KLOC (+ 250 ) N/A
Bootloader
5236 55KB
Security Monitor
9287 97KB
Mini-SM
2301 5.6KB
Linux
$\sim 20$ MLOC 157MB
SM Kernel Mod.
507 505KB
TCB Software
14523 152KB
Table 1. Software Component Size and TCB Breakdown. Mini-SM is a subset of the SM code. Lines in red correspond to elements excluded from the TCB.
Number of lines of code (LOC) and binary size of Citadel’s components are presented in Table 1. Due to their limited functions, the SM and the Bootloader can be kept small and the total software Trusted Code Base (TCB) is less than 15K LOC. This helps to minimize potential bugs and reduce the software attack surface. It can also help to make future formal verification efforts tractable. Final TCB size needs to include the enclave’s binary, which depends on the application.
6.3. Evaluating the End-to-End Platform
MicroPython Runtime We do a minimal port of MicroPython (microphython, ) inside one of our enclaves. We start with the bare-metal ARM implementation. We modify the configuration files and the linker script so our binary can run in the virtual memory layout provided by our enclave. We only need to implement a few functions by hand. For instance, we redirect the console to print to shared memory and need to write a bit of assembly for the setjmp and longjmp functions used for exception handling. We also add a cryptographic library, so the enclave can sign the console output using its own private keys recovered through the remote attestation mechanism. A hash of the standard output is extended on each character print and signed on the enclave exit. We cross-compile MicroPython using the RISC-V tool-chain, build without standard library, and generate position independent code. In total, we need to manually modify 128 LOC for the port and add the 5K LOC cryptographic library. The final compiled binary measures 180KB. With the enclave’s code added, the TCB is now 65,174 LOC, most of it (70%) being MicroPython.
Figure 6. MicroPython Enclave Execution Breakdown.
We use a simple primality test as a benchmark where a number is found to be prime if none of the numbers smaller than it can divide it. We test the primality of an arbitrary small prime number: $19,531$. The tests take an average of $4.1\times 10^{7}$ cycles to execute, with 51% of the time spent within the Python program. This means the enclave and attestation mechanism add 94% overhead for this small computation. We show a visual breakdown of the running time in Figure 6. We note that most of this overhead comes from hashing enclave code (25% of the total running time) but also generating the enclave keys and signing the certificate (10% of the running time). These two numbers are especially high since this code is run non-speculatively inside the SM. On another hand, hashing and signing the output inside of enclave takes a negligible amount of time (less than 1% of the total running time). Because this cost is fixed, it could definitely be amortized for bigger applications.
Securing a Cryptographic Library We write a wrapper around an ED25519 RISC-V library (ed25519github, ). The wrapper makes it possible to keep the keys used by the library inside of the enclave and for the library functions to be accessed in a Remote Procedure Call style (RPC). We add a queue to shared memory, so untrusted applications can send requests to the library to generate keys (that will stay in enclave memory), or to sign messages using previously generated keys. In total, we need to implement 334 LOC of C, which is to be compared to the 4,514 LOC of the crypto library (7% overhead). With the enclave’s code added, the total TCB is now 19,231 LOC. The final crypto-enclave binary is 71KB.
We compare performance between three scenarios. In each of these scenarios, we measure the number of cycles required to sign randomly generated messages. We split 1MB of randomly generated data into $\approx 3000$ packets of various lengths following the arbitrary Internet Mix distribution (imix, ). In the first setup, which corresponds to our insecure baseline, we do not use any enclave as we directly link the cryptographic library to our user-level application. In the second scenario, called “multi-core”, the enclave constantly runs on a dedicated core and programs running on other cores can forward requests through the shared memory mechanism. In the third scenario, termed “re-entry”, the enclave is re-entered every single time an application calls the cryptographic library on the same core as that of the application. We see very little difference (2.9%) between the first two setups, and a more important 27.6% slowdown between the first and third scenario. Indeed, our cryptographic algorithm is compute-heavy, making the communication overhead negligible in the second scenario. Nevertheless, context-switching in and out of an enclave is expensive as it requires flushing the microarchitectural state, which makes the slowdown more important in the re-entry scenario. These two secure solutions highlight an interesting tradeoff between resource occupation (one core and a DRAM region vs. only a DRAM region) and performance.
Private Inference We evaluate the following scenario where a remote user wants to do perform some private inference task on a neural network evaluated on an untrusted server. We pace a simple neural network model inside of one of our enclaves. We use onnx2c (onnx2cgithub, ) to convert a ONNX file describing a handwritten digit recognition model (trained on MNIST) to standard C code and run private inference inside of our enclave. The user uses our remote attestation mechanism to verify the enclave identity and establish a shared secret with it. The user is then able to send 1024 encrypted requests to the enclave that are passed along to the enclave by the untrusted OS. Upon reception, the enclave decrypts the ciphertext and performs the inference task on the private input. It then sends back an encrypted output to the remote user. In total, we need to implement 227 lines of straightforward C, which is to be compared to the 1536 LOC of the neural network and the 410 LOC of the AES library and the 4386 LOC of ED25519 we now need to include inside of the enclave for encryption and key exchange. With the enclave code added, the TCB is now 21K LOC and the final enclave binary measures 54KB.
We now compare the performance between our setup and an insecure baseline where the model is directly evaluated outside of an enclave and where the LLC is not partitioned. We do consider that incoming traffic from the remote user is encrypted and needs to be decrypted first. In both setups, runtime is split 55% and 45% between AES operations and inference. We see negligible (0.76%) overhead for running a model inside of the enclave, with all of it coming from the necessity to copy the ciphertext in enclave private memory before performing decryption. This is not surprising as the rest of the computation happens in enclave private memory and should not suffer from any slowdown. This doesn’t account for the time required to set up the enclave, which is a one-time cost but represents 23.3% of the cost of performing 1024 inference.
6.4. Evaluating Secure Shared Memory
We first demonstrate that naive public memory sharing is vulnerable to transient execution attacks by mounting an attack on an insecure version of Citadel where speculation on shared memory is allowed. The attack uses the LLC as a side channel and use a specter gadget located in enclave code to leak arbitrary enclave memory to untrusted code ruining on another core. Anecdotally, we were not able to successfully mount the attack on our secure prototype.
Table 2. Normalized runtime for signing and hashing three thousand random messages sent by the OS to an enclave. Runtime is normalized to the insecure baseline where the library is directly linked to the testbench program. We compare Citadel with two other configurations where the library code is not annotated for burst accesses to shared memory (- Burst) or all annotated (++ Burst).
[…]
We then evaluate the perfromance overhead of our comunication mechanism. We use a similar setup as Section 6 but only measure time spent computing the cryptographic function. The enclave performs $3000$ cryptographic operations (signature or cryptographic hash) over random inputs provided by the OS. We compare three setups. First, our final design with burst mode enabled around shared memory burst sequences (Citadel). Then Citadel without annotating code for shared memory burst sequences (- Burst). Finally, another (insecure) configuration where we annotate all the code to be run in burst mode (++ Burst). Our results are presented in Table 6.4. We report the runtime normalized to the insecure baseline (directly linked library and no LLC partitioning). We observe an overhead of 2.7% on the signature benchmark and 46.6% for the hashing benchmark. This is because the hashing application accesses shared memory much more frequently compared to the signing application that is more computationally bounded. The hashing function is more affected by the loss of the L1 cache. This difference between the benchmarks is also highlighted by our other results, where not annotating the code for burst shared accesses has an $397.3\%$ overhead for the hash application and a $19.4\%$ for the signing applicaiton. These results also demonstrate the efficiency of our burst mechanism when used properly.
We also investigate the performance of using this mode on the whole enclave code instead of the limited code segment when the enclave communicates. On our benchmarks, we observe that the performance overhead goes up to $65.4\%$ for the signing benchmark and $395.9\%$ for the hashing benchmark. We investigate the origin of this overhead and realize that miss speculation due to the deactivation of the RAS generates a lot of random load that pollutes the TLB, triggers many page walks and terribly slows down the virtual memory subsystem. This is due to many small functions being linked one after the other in the binary. Some of these functions start with loads using generic registers as addresses (for instance when accessing a pointer passed as an argument to the function). When encountering a ret instruction, the processor will first speculate that the following instructions should be executed. In this speculative window, it will start running the instructions of the function linked right after and perform loads or stores to random addresses from registers containing garbage. We confirm this hypothesis by instrumenting the binary and observing that when the RAS is reactivated (but not other predictors), the number of TLB misses, page walks and the overall execution time go down significantly. As a result, we recommend a hybrid approach where the burst mode is only used when the enclave wants to communicate.
Table 2. Normalized runtime for signing and hashing three thousand random messages sent by the OS to an enclave. Runtime is normalized to the insecure baseline where the library is directly linked to the testbench program. We compare Citadel with two other configurations where the library code is not annotated for burst accesses to shared memory (- Burst) or all annotated (++ Burst).

### v4

@@
-Jules Drean MIT CSAIL Miguel Gomez-Garcia MIT CSAIL Fisher Jepsen MIT CSAIL Thomas Bourgeat EPFL  and  Srinivas Devadas MIT CSAIL
+Citadel: Real-World Hardware-Software Contracts for Secure Enclaves Through Microarchitectural Isolation and Controlled Speculation
 Abstract.
-Enclaves or Trusted Execution Environments are trusted-hardware primitives that make it possible to isolate and protect a sensitive program from an untrusted operating system.
-Unfortunately, almost all existing enclave platforms are vulnerable to microarchitectural side channels and transient execution attacks, and the one academic proposal that is not does not allow programs to interact with the outside world.
-We present Citadel, to our knowledge, the first enclave platform with microarchitectural isolation to run realistic secure programs on a speculative out-of-order multicore processor.
-We show how to leverage hardware/software co-design to enable shared memory between an enclave and an untrusted operating system while preventing speculative transmitters between the enclave and a potential adversary.
-We then evaluate our secure baseline and present further mechanisms to achieve reasonable performance for out-of-the-box programs.
-Our multicore processor runs on an FPGA and boots untrusted Linux from which users can securely launch and interact with enclaves.
-To demonstrate our platform capabilities, we run a private inference enclave that embed a small neural network trained on MNIST.
-A remote user can remotely attest the enclave integrity, perform key exchange and send encrypted input for secure evaluation.
-We open-source our end-to-end hardware and software infrastructure, hoping to spark more research and bridge the gap between conceptual proposals and FPGA prototypes.
+Hardware isolation primitives such as secure enclaves aim to protect sensitive programs, but remain vulnerable to transient execution attacks.
+Complete microarchitectural isolation is not a satisfactory defense mechanism as it leaves out public shared memory, critical for usability and application performance.
+Conversely, hardware-software co-designs for secure speculation can counter these attacks but are not yet practical, since they make assumptions on the speculation modes, the exposed microarchitectural state, and the software, which are all hard to support for the entire software stack.
+This paper advocates for processors to incorporate microarchitectural isolation primitives and mechanisms for controlled speculation, enabling different execution modes.
+These modes can restrict what is exposed to an attacker, effectively balancing performance and program-analysis complexity.
+We introduce two mechanisms to securely share memory between an enclave and an untrusted OS in an out-of-order processor.
+We show that our two modes are complementary, achieving speculative non-interference with a reasonable performance impact, while requiring minimal code annotation and simple program analysis doable by hand.
+Our prototype, Citadel, is a multicore processor running on an FPGA, booting untrusted Linux, and supporting comprehensive enclave capabilities, such as shared memory, and remote attestation.
+To our knowledge, Citadel is the first end-to-end enclave platform to run secure applications, such as cryptographic libraries or small private inference workloads, on a speculative out-of-order multicore processor while protecting against a significant class of side-channel attacks.
 1. Introduction
-Secure Enclaves or Trusted Execution Environments (TEEs) (costan2016sanctum, ; aegis, ; ferraiuolo2017komodo, ; lee2020keystone, ; bahmani2021cure, ; feng2021scalable, ), are trusted-hardware primitives used to protect the execution of sensitive programs.
-They make it possible to run sensitive computations in isolation with strengthen privacy and integrity properties.
+Hardware isolation primitives such as Secure Enclaves or Trusted Execution Environments (TEEs) (costan2016sanctum, ; aegis, ; ferraiuolo2017komodo, ; lee2020keystone, ; bahmani2021cure, ; feng2021scalable, ) are used to run sensitive computations in isolation with strengthen privacy and integrity properties, even in the presence of a malicious Operating System (OS) or hypervisor (see Figure 1).
 Academic TEE proposals date back twenty years to XOM (xom, ) and AEGIS (aegis, ) and industry implementations include ARM TrustZone (trustzone04, ; pinto2019demystifying, ), Intel SGX (mckeen2013sgx, ; costan2016intel, ) and AMD SEV (sev2020strengthening, ).
-Unfortunately, these kinds of platforms have been shown to be vulnerable to microarchitectural side channels and transient execution attacks (chen2019sgxpectre, ; koruyeh2018spectre, ; ryan2019hardware, ; van2020lvi, ).
-In particular, attacks such as those in (van2018foreshadow, ; van2020sgaxe, ) were able to not only extract sensitive data from Intel SGX enclaves, but also managed to break the security of the remote attestation mechanism.
+Unfortunately, these kinds of platforms have been shown to be vulnerable to microarchitectural side channels and transient execution attacks (chen2019sgxpectre, ; koruyeh2018spectre, ; ryan2019hardware, ; van2020lvi, ; van2018foreshadow, ; van2020sgaxe, ).
 One academic proposal (bourgeat2019mi6, ) has looked at securing enclaves in the context of microarchitectural side channels and transient execution attacks.
-They aim for a strong security guarantee where a potential attacker can only observe the enclave’s output.
-Unfortunately, this extreme isolation strongly restricts the class of programs that can be run, as no communication with the outside world is allowed except when entering or exiting the enclave.
+They aim for total microarchitectural isolation between a (victim) enclave and the other security domains collocated on the same machine, leaving out shared memory, a commonplace mechanism at the center of many software and applications.
 Figure 1. Citadel’s Computing Stack
-We are not the first to observe that restrictive enclave execution models limit the programmability and the performance of enclave programs(van2023cheri, ; yu2022elasticlave, ; li2021confidential, ; lee2022cerberus, ; park2020nested, ; steinegger2021servas, ) but those work ignore security challenges introduced by microarchitectural side-channels and transient execution attacks.
-When considering those challenges, even basic communication primitives like public shared memory with the operating system are not straightforward to implement efficiently and securely.
-In this paper, we try to fill the gap between platform that have strong security but restrict enclave programs to the extreme and platforms that allow better enclave execution models but ignore side channels and transient execution attacks.
-We first define our relaxed security goal: an attacker should only be able to observe the transcript of non-speculative accesses to a shared memory region.
-We then craft requirements for the hardware and the software so that, together, they can prevent secrets from leaking through speculative access to shared memory.
-In practice, the hardware will isolate private microarchitectural state and restrict the victim program speculation model to help it achieve the just described security guarantee.
-Intuitively, the more restrictive the speculation model, the lighter the requirements on the program, but the least performance we get.
-In our case, we first consider a very conservative speculation model where no speculative shared accesses are allowed.
-This effectively prevents the existence of any speculative transmitters between the victim and the attacker.
-We describe how to enable shared memory securely for the specific open-source out-of-order design we used, while disabling speculative shared accesses.
-We also make sure to not enable any new speculative transmitters by bypassing the L1 D cache, and to not break address translation by tagging the TLB’s and the Address Translation Cache’s entries.
-Finally, we give a brief overview of the software infrastructure we have implemented in order to boot Linux and run enclaves from user programs on our FPGA prototype.
-We use this secure baseline to evaluate the performance of the kind of programs that one could be interested in putting in an enclave.
-Out-of-the-box, we obtain an underwhelming $120.6\%$ performance overhead for a cryptographic library.
-We identify and analyze the performance bottlenecks to design a series of hardware-software mechanisms to improve both enclaves’ and unprotected processes’ performance.
-More precisely, we develop a reconfigurable Last Level Cache (LLC) partitioning scheme, a performance necessity for our small LLC.
-This reduces our overhead from $120.6\%$ to $19.4\%$.
-We also introduce software controls to enable a slightly less conservative enclave speculation model, where pipelined accesses to shared memory are allowed but predictors are hardwired to $pc+4$.
-The efficient way to use this mode is on code segments with burst accesses to shared memory.
-Under this less conservative speculation model, the programmer needs to make sure that straight-line speculation ($pc+4$) within these segments will not leak any secrets.
-This mechanism reduces overheads from $19.4\%$ to $2.7\%$ on our cryptographic benchmark.
+On the other hand, mechanisms for secure speculation (ainsworth2020muontrap, ; cheang2019formal, ; canella2019systematic, ; barber2019specshield, ; fustos2019spectreguard, ; yan2018invisispec, ; yu2019speculative, ; saileshwar2019cleanupspec, ; weisse2019nda, ; khasawneh2019safespec, ; guarnieri2020spectector, ; guarnieri2021hardware, ; mosier2022axiomatic, ) can thwart these types of attacks while enabling shared resources but are not yet practical.
+Solutions that are purely hardware are complex: none has been implemented beyond simulators, and attacks have been found for the majority of schemes.
+Hardware-software co-design approaches simplify that problem by only providing guarantees for some programs, but they make strong assumptions on the type of speculative execution, the exposed microarchitectural state, and the software, which are hard to support for the entire software stack.
+We propose that processors expose microarchitectural isolation primitives and mechanisms for controlled speculation.
+Indeed, we believe these two approaches go hand-in-hand as isolation makes it possible to constrain the problem of secure-speculation enough so hardware-software co-design solutions only require minimum hardware changes and a tractable program analysis component.
+To demonstrate this idea, we present two mechanisms that securely enable shared memory between an enclave and an untrusted OS on an out-of-order processor.
+Together, the enclave program and the hardware achieve Speculative Non-Interference (SNI) (guarnieri2020spectector, ; guarnieri2021hardware, ; behnia2021speculative, ; fabian2022automatic, ; yu2019speculative, ; yu2019speculative2, ) i.e., a victim program running in an enclave does not leak more secret information when running speculatively using our mechanisms than when running non-speculatively (one-instruction-at-a-time and in-order).
+For the rest of this paper, we will refer to non-speculative execution as sequential execution.
+Our security argument leverages hardware software contracts for security (guarnieri2021hardware, ; mosier2022axiomatic, ).
+Through microarchitectural isolation, the attacker observations are restricted to the addresses of victim accesses to shared memory.
+The hardware can then enforce some controlled speculative execution mode for the enclaves, such that an analyzed program can be run securely with regard to that mode.
+Intuitively, the more restrictive the execution mode, the lighter the requirements on the program, but the least performance we get.
+Concretely, we first propose and build SpecSafe, a hardware mechanism that enforces an execution mode for the enclave where most instructions can be run speculatively, but no speculative shared accesses are allowed.
+This effectively prevents the existence of any speculative transmitters between the victim and the attacker without requiring any program modification or analysis.
+Unfortunately, we show it has limited performance on code that perform back-to-back accesses to shared memory.
+As a result, we present Burst Mode, a second enclave execution mode.
+We introduce software controls to enable Burst Mode where pipelined accesses to shared memory are allowed, but all the predictors of the executing core are hardwired to $pc+4$.
+This speculation model is usually referred to as “straight-line speculation” (straightlinespecARM, ) and we explain under which conditions a program can use it securely.
+We show that program analysis is tractable by hand on small straight-line code segments with back-to-back accesses to shared memory (e.g., a memcopy function).
+Our two execution modes are complementary, achieving SNI with a reasonable performance impact, while requiring minimal code annotation and simple program analysis doable by hand.
+To our knowledge, we are the first to prototype this type of security contract on a full platform, i.e., implemented on a processor running on FPGA booting Linux.
+While building this prototype, we uncovered a few implementation subtleties and potential security issues that had been overlooked in the past, and we prototyped and run a simple-to-implement and effective Last Level Cache (LLC) isolation technique.
+Those side quests, though more implementation-centered, would be especially important to potential industrial deployment.
 Our final prototype is a modified out-of-order RISC-V 2-core processor synthesized to run on an AWS F1 FPGA.
-It fully implements all our security and dynamic resizing mechanisms, boots (untrusted) Linux and makes it possible for user-level applications to launch and safely communicate with secure enclaves executing real applications such as a lightweight Python runtime, a cryptographic library or a private inference application running a small neural network.
-We also implement an end-to-end attestation mechanism, making it possible for an enclave to identify itself independently of where the OS placed it in physical memory, enabling enclaves to act as a trusted third party.
-We implement and make all the pieces of software and modified hardware available, such as our secure bootloader, our security monitor and a Linux kernel module to interact with the security monitor.
+It fully implements all our security and performance mechanisms at the microarchitectural level, along with an end-to-end attestation mechanism.
+It boots (untrusted) Linux and user-level applications can safely launch and communicate with secure enclaves, executing applications such as a lightweight Python runtime, a cryptographic library or a private inference application running a small neural network.
+We implement and make all the pieces of software and modified hardware available, such as our modified processor, our secure bootloader, our security monitor and a Linux kernel module to interact with the security monitor.
 We summarize the different elements implemented for this paper in Figure 1.
-This makes our platform a unique environment for research in the domain of microarchitectural side channels, transient execution attacks and secure enclaves.
-We hope researchers will take ownership of the platform, attack it, modify it to prototype and test complex hardware-software mechanisms and research ideas.
+Citadel offers a unique environment for testing complex hardware-software mechanisms, making it ideal for research into microarchitectural side channels, transient execution attacks, and secure enclaves.
 To summarize, the main contributions of the paper are:
-  • To our knowledge, we show the first end-to-end enclave platform running securely on an FPGA implementation of a speculative out-of-order processor.
-  • We show how to build a shared memory mechanism while guaranteeing an attacker will only observe non-speculative accesses to shared memory.
-  • Guided by early performance evaluation, we propose a series of hardware-software mechanisms to make the platform more configurable and improve both enclaves’ and unprotected processes’ performance.
+  • To our knowledge, we show the first implementation leveraging hardware-software contracts to guarantee well-defined security properties.
+  • We demonstrate how microarchitectural isolation goes hand-in-hand with controlled-speculation by presenting two mechanisms for public shared memory on enclaves running speculatively.
+  • To our knowledge, we show the first end-to-end enclave platform prototype running securely on an FPGA implementation of a speculative out-of-order machine.
+  • We show a hardware-software mechanism for cache partitioning that makes the platform more configurable and improves both enclaves’ and unprotected processes’ performance.
+  • Through our case study, we provide important takeaways for commercial enclave offerings.
 We open-source our hardware, software and debugging infrastructure at https://github.com/mit-enclaves/Citadel.
-2. Motivation & Background
-We want to : 1) run programs that can interact with the outside world. 2) formalize the guarantees that we expect the hardware to provide 3) have good performance.
+2. Background
 2.1. Security Monitor
-In many enclave platforms, including Citadel, the Security Monitor (SM) (costan2016sanctum, ) (lebedev2019sanctorum, ) is a small (9K LOC), trusted piece of software running at a higher privilege mode than the hypervisor or the OS.
+In many enclave platforms, including Citadel, the Security Monitor (SM) (costan2016sanctum, ) (lebedev2019sanctorum, ) is a small (9K LOC in our prototype), trusted piece of software running at a higher privilege mode than the hypervisor or the OS (see Figure 1).
 Its role is to link low-level invariants exposed by the hardware (e.g., if a specific memory access is authorized), invoking the right microarchitectural cleanup routine, and the high-level security policies defined by the platform (e.g., which enclave is currently executing on which core and which memory region does it own).
@@
 Cache side channels are the easiest channels to exploit, and so have been extensively studied, but any piece of shared microarchitectural state can theoretically be used as well.
-To summarize the observations that could be made through side-channels, the community formalizes side-channels as a set of observations that an attacker could make about the microarchitectural state.
-For example, cache side-channels on unprotected machines typically lead to an observation model that includes the sequence of addresses of memory operations (sometimes with their timestamps) as part of the observation.
-2.3. Transient Execution Attacks and Execution Models
+To summarize the observations that could be made through side-channels, the community formalizes side-channels as a set of observations that an attacker can make about the microarchitectural state.
+For example, cache side-channels on unprotected machines typically lead to an observation mode that includes the sequence of addresses of memory operations (sometimes with their timestamps).
+2.3. Transient Execution Attacks
 Speculative or transient execution is one of the key performance features of modern processors.
@@
 Some microarchitectural data structures that orchestrate speculation (e.g., branch predictors) can be trained so the processor will execute code snippets speculatively, in ways that are incorrect or on data that should not be accessed during normal execution.
-Concretely, instructions that were not part of the program normal execution path can be executed out-of-thin-air, speculatively, by the processor.
+Concretely, instructions that were not part of the normal program execution path can be executed out-of-thin-air, speculatively, by the processor.
 An example is the execution of code speculatively before the resolution of a preceding security check, such as a memory access bound.
@@
 Nonetheless, they will still change the microarchitectural state of the machine.
-For example, a speculative load to some data will not affect the values of the architectural registers but will load the data in the memory hierarchy and possibly change the state of caches.
+For example, a speculative load to some data will not affect the values of the architectural registers, but will load the data in the memory hierarchy and possibly change the state of caches.
 This means that a program that manages sensitive data can be coerced into leaking private data speculatively through a side channel.
@@
 This type of attack, called Spectre, was first introduced in (Kocher2018spectre, ), and many variants have since been demonstrated.
+2.4. Hardware Software Contract for Security
+A recent line of work (guarnieri2021hardware, ; mosier2022axiomatic, ) has been trying to define hardware software contracts for security, also called security contracts.
+As said in (guarnieri2021hardware, ), “the purpose of a contract is to split the responsibilities for preventing side channels between software and hardware”.
+An initial observation is that not all programs pose the same level of problem or difficulty in mitigation for all microarchitectural side-channels.
+This line of work tries to identify and describe classes of programs for which it is possible to implement performant microarchitecture securely, i.e., the programs should not leak more secrets through the microarchitecture implementation, than what it would directly leak in a non-speculative or sequential (one-instruction-at-a-time and in-order) Instruction Set Architecture model.
+In this work, we reuse the framework from (guarnieri2021hardware, ) while abstracting away some of its formalism.
+We point out that if we make some simple assumptions (that we check) on the parts of the enclave programs that are in charge of interacting with shared memory, we can run those parts of the enclave in a controlled speculation mode, preserving performance while ensuring strong security guarantees.
 3. Defining Security
-To achieve security, we rely on both the hardware to enforce some isolation and on the fact the software verifies some requirements: we do not protect all-software out-of-the-box.
-Intuitively, for a given security property, if the isolation enforced by the hardware is stronger, then the requirements on the software are weaker and we can enforce the security property for more programs.
-Previous Work Previous work looking at enclaves in a context of transient execution attacks (bourgeat2019mi6, ) aims at limiting the attacker’s view to the enclave final output.
-Completion time was considered outside of the threat model and could be independently handled using traffic shaping techniques (padding) techniques.
-Unfortunately, this security property is too restrictive as it only considers a limited class of programs that do not communicate with the outside world.
-Public Shared Memory We define a weaker security property in order to consider a more important class of programs to run securely.
-Our goal is for the system to guarantee that an attacker should only be able to observe the transcript of non-speculative accesses to a shared memory region.
-In particular, we don’t want an attacker to be able to observe any speculative accesses to shared memory from the victim nor any other microarchitectural state (accesses to private memory, branch predictors, TLBs etc…).
-However, similarly to previous work, we do not consider the timing of each access is observable by the attacker, i.e., there are no timestamps for the non-speculative shared memory accesses.
-We consider that leakage through the timing of shared memory accesses should be addressed using orthogonal traffic-shaping techniques such as (deutsch2022dagguise, ; wang2014timing, ; shafiee2015avoiding, ; zhou2017camouflage, ) and is left for future work.
-Hardware & Software Requirements We now define what kind of requirements the hardware will needs to enforce and how the hardware restricts the victim program speculation model to help it achieve the just described security guarantee.
-1) the hardware should properly isolate private microarchitectural state.
-2) together, the hardware and the software need to prevent secrets from leaking through speculative accesses to shared memory.
-Intuitively, the more restrictive the speculation model of shared memory accesses enforced by the hardware, the lighter the conditions on the program to finally achieve security, but the least performance we get.
-We first experiment with a restrictive speculative model where the victim program is not able to perform speculative accesses to shared memory.
-This effectively prevents the existence of any speculative transmitters between the victim and the attacker.
-The program only needs to make sure that the explicit shared memory accesses don’t leak information about the program secrets.
-Any constant-time cryptographic code satisfies these restrictions.
-Unfortunately this restrictive model gives us poor performance.
-We experiment with a second speculative model called Burst mode where speculation on shared memory accesses is allowed, but all the predictors are deactivated and hard-coded to $pc+4$.
-The rationale is to allow pipelined accesses to shared memory to improve performance.
-Burst mode is intended to be only activated for code segments with straight-lined, burst accesses to shared memory.
-Under this less conservative speculation model, the software requirements are more important and the programs needs to verify that straight-line speculation ($pc+4$) within these segments will not leak any secrets.
-Outside of the said segments, we use the first speculation model.
+Previous Work Previous work looking at enclaves in a context of transient execution attacks (bourgeat2019mi6, ) enforces Strong Microarchitectural Isolation (SMI) by completely isolating security domains at the microarchitectural level.
+See Section 4.1 on hardware isolation.
+Unfortunately, this isolation forbids any shared architectural state between an enclave and the OS, disallowing critical functionalities such as shared memory.
+Any interaction between the enclave and the OS has to go through the SM and requires several (expensive) context switches, microarchitectural flushes and data copies from one domain to another.
+In (bourgeat2019mi6, ) enclave completion time is considered outside of the threat model.
+Generally, completion time can and should be handled independently using traffic shaping (padding) techniques.
+Secure Shared Memory We relax SMI to allow some shared memory between the enclave and the OS.
+Nevertheless, to prevent transient execution attacks, we require that the enclave software and the hardware together achieve Speculative Non-Interference (SNI) (guarnieri2020spectector, ; guarnieri2021hardware, ; behnia2021speculative, ; fabian2022automatic, ; yu2019speculative, ; yu2019speculative2, ), i.e., that a victim program running in an enclave should not leak more secrets when running speculatively using our mechanisms than if run with non-speculatively (sequential execution).
+However, similar to previous work, we do not consider the timing of microarchitectural events to be visible to the attacker.
+In summary, no timestamps are associated with the observed shared memory accesses and the associated microarchitectural events; only the traces of the corresponding addresses are exposed.
+We consider that leakage through the timing of shared memory accesses is similar to leakage through completion time, and it should be addressed using orthogonal traffic-shaping techniques such as (deutsch2022dagguise, ; wang2014timing, ; shafiee2015avoiding, ; zhou2017camouflage, ).
+This is left for future work.
 3.1. Attacker Capabilities
 We now complete our threat model by describing the attacker’s capabilities.
-The attacker is collocated on the same machine as the victim and can mount any software-based attack, including microarchitectural side channels and transient execution attacks.
-The attacker has compromised the majority of the software stack, including the operating system (OS) except for a limited trusted code base (TCB) including the secure bootloader, the SM (together 15K lines of code (LOC)) and the enclave.
+The attacker is collocated on the same machine as the victim enclave and can mount any software-based attack, including microarchitectural side channel and transient execution attacks.
+The attacker has compromised the majority of the software stack, including other malicious enclaves and the operating system (OS), except for a limited trusted code base (TCB) including the secure bootloader, the SM (together 15K lines of code (LOC)) and the victim program running inside of the enclave.
+The attacker also has control over the contents of shared memory.
 For the rest of this paper, we refer to any code outside of the TCB as the OS.
@@
 As highlighted by our security definition, we are worried about enclave privacy and do not consider denial of service.
-4. Citadel 1.0: A Secure Baseline
-Even though defense mechanisms often propose to ”not speculate” when executing some instructions that can dirty the microarchitectural state, implementing such policies is not completely straightforward.
-In this section, we describe how to securely enable shared memory for our specific out-of-order open-source design while disabling speculative shared accesses.
-We also make sure to not enable any new speculative transmitters by bypassing the L1 D cache, and to not break address translation by tagging the TLB’s and the Address Translation Cache’s entries.
-Finally, we give a brief overview of the software infrastructure we have implemented in order to boot Linux and run enclaves from user programs.
-4.1. Hardware Isolation
-Memory Isolation Sanctum (costan2016sanctum, ) already presents hardware mechanisms for memory isolation in the context of an untrusted OS.
-Like RISC-V Physical Memory Protection (PMP) (waterman2015risc, ), Sanctum introduces a mechanism to partition physical memory and prevent a malicious OS from accessing another security domain’s private memory regions, or different enclaves from accessing each other’s private memory.
+We also do not protect enclave programs that explicitly leak their own secrets, including through shared memory.
+4. Secure Shared Memory
+In this section, we first describe our baseline microarchitecture, and how it enforces strong isolation between security domains with shared memory disabled.
+We then describe how to enable shared memory, how it opens the door to Spectre-like attacks, and why existing mechanisms are insufficient to address them.
+We then describe SafeSpec and Burst Mode, our two complementary mechanisms to securely and efficiently enable shared memory on speculative microarchitectures.
+4.1. Strong Microarchitectural Isolation
+Figure 2. Resource Partitioning Between Security Domains
+Figure 3. Context Switching Between Security Domains
+Memory Isolation Sanctum (costan2016sanctum, ) introduces hardware mechanisms for memory isolation in the context of an untrusted OS.
+Like RISC-V Physical Memory Protection (PMP) (waterman2015risc, ), Sanctum introduces a mechanism to partition physical memory and prevent unprivileged security domains (OS or enclave) from accessing the private memory regions of another security domain (OS, enclave or SM).
 At its core, the mechanism uses two 64-bit registers per core (one for trusted memory and one for untrusted/shared memory) as memory bitmaps.
-Each bit represents a fixed-size physical memory region (typically 32MB) and functions as an access-rights flag for the currently-running program.
-Citadel borrows this mechanism for its simplicity and enforces these checks on every memory access, including speculative accesses and for accesses performed during page walks.
-Microarchitectural Isolation Following MI6 (bourgeat2019mi6, ) methodology we partition all spatially shared microarchitectural state (LLC, MSHRs…) and flush temporally shared state on context switch (all cores’ pipeline state, L1 caches, TLBs…).
-We are not vulnerable the DRAM controller side channels as we inherit MI6’s constant time DRAM padding mechanism.
-Finally, our processor does not support hyper-threading and we are not vulnerable to related side channels such as port contention.
+Each bit represents a fixed-size physical memory region (typically 32MB) and functions as an access-rights flag for the currently-running security domain.
+Citadel borrows this mechanism for its simplicity and enforces these checks on every memory access, including speculative accesses and on accesses performed during page walks.
+Memory isolation is illustrated in Figure 2, where the coloring highlights that only currently running security domains can access corresponding DRAM memory regions.
+Microarchitectural Isolation Following MI6 (bourgeat2019mi6, ), we enforce strong microarchitectural isolation through partitioning, flushing and by disabling speculation in the SM.
+First, we partition all spatially-shared microarchitecture including the LLC, the LLC pipeline and Miss Status Handling Registers (MSHRs) and other resources as illustrated in Figure 2.
+In particular, the LLC is statically partitioned into 64 regions to match the 64 DRAM memory regions – we will improve on this scheme in Section 6.
+We pad the timing of DRAM requests to a constant, which eliminates DRAM controller side channels.
+When context switching between security domains, we leverage interrupt logic to interpose the SM and enforce the flushing of temporally-shared microarchitecture such as all cores’ pipeline state, L1 caches, TLBs, Translation Caches and MSHRs (see Figure 3).
+The processor does not support hyper-threading and therefore is not vulnerable to related side channels such as port contention.
+The processor does not include a prefetcher so we are not susceptible to Spectre variants exploiting them.
+Finally, Citadel is not susceptible to Meltdown-style attacks, as the processor does not speculate across privilege levels (switching to a different privilege mode flushes the re-order buffer, making sure it never contains two instructions from different privilege levels).
+The exhaustive list of side channels and defense mechanisms covered by MI6 can be found in (bourgeat2019mi6, ).
+Disabling Speculation in the SM The SM executes in machine mode (see Figure 3) which gives it access to all physical memory, making it vulnerable to Spectre-like attacks.
+Following (bourgeat2019mi6, ), we disable speculation in the SM to prevent these attacks.
 4.2. Enabling Shared Memory
-Figure 2. The Dual Page Table Mechanism
-Sanctum also uses a dual-page-table mechanism (see Figure 2) to differentiate between trusted and untrusted memory based on virtual addresses, which we also borrow in Citadel.
-Two new registers (equivalent to base and bounds) are used to define the trusted virtual memory range (i.e.
-the enclave’s virtual memory).
+Figure 4. The Dual Page Table Mechanism + Restricting Speculative Access to Shared Memory
+Sanctum, a non-speculative design, proposed a dual-page-table mechanism (see Figure 4) to differentiate between trusted and untrusted memory based on virtual addresses and enable shared memory.
+We adopt this dual-page-table mechanism in Citadel.
+Two new registers (equivalent to base and bounds) are used to define the trusted virtual memory range (i.e., the enclave’s virtual memory).
 The SM enforces the invariant that the physical memory mapped as trusted is owned by the corresponding enclave.
@@
 This mechanism is especially handy to eliminate speculative transmitters, as there is no need to perform any speculative address translation (which could leak information) to know whether a memory access is trusted or not.
-4.3. Restricting Speculative Access to Shared Memory
+4.3. Attacks and Reusing Existing Mechanisms
+Shared memory implies that security domains now share some microarchitectural state, specifically a shared LLC partition.
+This effectively breaks Strong Microarchitectural Isolation and opens the door to LLC side channels and Spectre-like attacks.
+Anecdotally, we have successfully mounted such an attack against an enclave on our insecure baseline running on an FPGA.
+We were not able to successfully mount the attack on our final secure design.
+To thwart these attacks, we need to achieve SNI (see Section 3) i.e., prevent speculative accesses to shared memory from leaking enclave secrets, for instance, by preventing speculative accesses to shared memory altogether (see Figure 4).
+Nevertheless, building such a mechanism is not trivial.
+For instance, existing mechanisms that prevent some speculative memory accesses (MMIO accesses or non-cacheable memory (schwarz2020context, )) cannot be repurposed.
+Indeed, they rely on physical addresses or page table entries to identify memory regions, implying that they allow some speculative memory accesses to page tables, already leaking part of the address.
+Instead, we propose to design new mechanisms for speculation-safe shared memory to balance security and performance.
+4.4. SpecSafe: Speculatively Safe Shared Memory
 In the baseline out-of-order processor we are hardening, memory instructions are decoded, renamed and then simultaneously queued in the reorder buffer (ROB) and the memory reservation station (MemRS).
@@
 These changes require the modification of MemRS and MemExePipeline.
-Figure 3. Changes to Memory Execution Pipeline. New elements are shown in blue.
-Modifying the Memory Execution Pipeline Upon entry into the MemExePipeline, checks are performed on memory instructions to decide if the memory operation is safe to translate and then issue, or not.
+Figure 5. Changes to the Memory Execution Pipeline to enable SpecSafe. New elements are shown in blue.
+Modifying the Memory Execution Pipeline Upon entry into the MemExePipeline, checks are performed on the memory instruction to decide if the memory operation is safe to translate and then issue, or not.
 It is important to perform these checks very early as the next steps in the MemExePipeline touch shared microarchitecture (through address translation and cache hierarchy for instance) potentially transmitting information to an attacker.
 A memory access is deemed safe if it accesses trusted memory OR if it is non-speculative.
-Because the enclave trusted memory range is defined in virtual memory, this check can be performed as soon as operands (i.e.
-virtual address) are available.
+Because the enclave trusted memory range is defined in virtual memory, this check can be performed as soon as operands (i.e., virtual addresses) are available.
 If an instruction tries to access untrusted or shared memory, we check that it is non-speculative by conservatively checking if it has reached the head of the ROB.
@@
 If it is not safe, the instruction is squashed and a signal is sent to the MemRS indicating that it should re-dispatch it once it reaches the head of the ROB.
-Modifying the Memory Reservation Station Our mechanism requires MemRS being able to re-dispatch some instructions once they become non-speculative: untrusted memory loads will indeed typically need to get dispatched twice.
+Modifying the Memory Reservation Station Our mechanism requires MemRS being able to re-dispatch some instructions, once they become non-speculative – untrusted memory loads will indeed typically need to get dispatched twice.
 Disabling the Mechanism for the OS Blocking speculation on memory access can have a significant performance impact.
-The SM makes sure that nonenclave programs (typically the OS and standard usermode programs) are still allowed to speculate for their memory accesses by disabling the trusted memory mechanism.
-4.4. Speculative Transmitters in the L1 D Cache
-The core-private part of the memory system, including the L1 D cache, is now shared between trusted memory accesses and untrusted or shared memory accesses.
+The SM makes sure that non-enclave programs (typically the OS and standard user mode programs) are still allowed to speculate when accessing their own memory by disabling the trusted memory mechanism.
+Note that untrusted code is still barred form (speculatively) accessing enclave or SM memory, thanks to memory isolation (see Section 4.1).
+4.5. Burst Mode: Pipelined Shared Memory Accesses
+Figure 6. Changes to the Fetch Pipeline to enable Burst Mode. New elements are shown in blue.
+We observe that SpecSafe performs poorly on code that performs back-to-back accesses to shared memory.
+This is mainly because shared memory accesses cannot be pipelined, as each instruction needs to wait to reach the head of the ROB to be sent.
+This serialization problem is exacerbated as we cannot use the L1-D cache for security reasons (see Section 6), increasing latency of responses.
+In some cases, this behavior is even exacerbated by the compiler, trying to optimize the code by placing all the load instructions to shared memory, one after the other.
+This is especially costly during burst accesses.
+The LLC supporting 16 outstanding requests and with a 10 cycle latency, means that 16 memory accesses should be pipelined to take a total of $10+16=26$ cycles.
+Without pipelining, they will now take $10\times 16=160$ cycles.
+As a result, we designed a complementary mode where we temporarily change the execution mode to allow shared memory burst accesses, but without enabling any type of attacker-controlled speculation.
+For instance, we don’t want our program to speculatively jump to a Spectre-like gadget and transmit secrets to our attacker.
+To avoid this, we restore speculative accesses to shared memory, but disable the use and training of all predictors, besides “pc+4”.
+Specifically, we modify the Fetch pipeline in order to disable the Branch Target Buffer (BTB), the Branch History Table (BHT) and the Return Address Stack (RAS) (see Figure 6).
+This is commonly referred to as straight-line speculation (straightlinespecARM, ; guarnieri2021hardware, ).
+Note that, for security, we still need to bypass the L1-D (see Section 6).
+Proper Usage The switch between SafeSpec and Burst mode is triggered by writing to a specific Control Status Register (CSRs).
+Note that CSR reads and writes cannot be performed speculatively, and the corresponding instructions effectively act as speculation barriers by flushing the processor pipeline.
+The right way to use this mode is on small code snippets that perform back-to-back accesses to shared memory, as shown in Figure 7 or Listing 2 (full code and assembly in supplementary material).
+The programmer or compiler only needs to identify these series of in-lined burst accesses and guard them with the CSR write instructions activating and deactivating burst mode, effectively isolating the code snippets.
+Modifying the execution mode also modifies the requirements on the software.
+In Figure 7, we highlight how the two mechanisms change the (speculative) control flow for the same assembly snippet from memcpy.
+The programmer now needs to ensure that the victim program does not leak secrets under straight-line speculation.
+More details and examples on how to verify a program’s compliance can be found in Section 5.
+We observe significant performance improvement when using Burst on our small code snippets.
+Interestingly enough, we also observe that (insecurely) activating burst mode on the whole enclave code leads to disastrous performance as it completely pollutes the TLB.
+More results and details are presented in Section 7.1.
+This highlights how Burst mode is complementary to SpecSafe and the importance of software-exposed controls for execution modes.
+5. Security Analysis
+Figure 7. Speculative control flow of memcpy assembly under our two mechanisms for secure shared memory
+Citadel builds upon MI6, which already enforces Strong Microarchitectural Isolation (SMI) through partitioning and flushing (See Section 4.1 and (bourgeat2019mi6, ) for details).
+As a result, our security analysis focuses on the novelties brought forward by Citadel, i.e., shared memory between an enclave and the OS.
+This change effectively opens side channels in the shared memory hierarchy through the shared LLC partitions.
+Note that any other side channels (e.g., branch predictors) are still closed per inherited MI6’s isolation mechanisms: partitioning and flushing.
+Furthermore, we only worry about the confidentiality of the enclave with regard to the OS here (see Section 3).
+As explained in Section 3, the high-level security guarantee we are trying to enforce is Speculative Non-Interference (guarnieri2020spectector, ) (SNI) between the attacker and the victim program.
+To formally reason about SNI and capture these new interactions, we need a different set of arguments than for SMI and to leverage the framework of hardware software contracts for security introduced in Section 2.4.
+Following this methodology, we first define the security contracts we will consider.
+5.1. Defining Security Contracts
+A security contract essentially defines what architectural state of the victim is considered exposed or public.
+The contract is defined using two building blocks: the observer mode, or what architectural state is exposed, and the execution mode or which paths of a given program are considered (speculatively) reachable and exposed.
+We define an observer mode that views the sequence of addresses of every (speculative) victim access to shared memory, and the data of non-speculative accesses, but no other victim architectural state (i.e., the enclave PC is not exposed).
+We also consider that the sequence does not contain any timing information (i.e., no timestamps, only data and addresses).
+We call this observation mode “shared memory” abbreviated to shm.
+We now describe the two different execution modes we consider for our contracts.
+The first execution mode we consider is the sequential mode (no speculation at all, i.e., one-instruction-at-a-time and in-order) or seq, which means that only non-speculative execution paths are considered reachable and addresses and data of shared memory accesses are exposed.
+We call this first contract seq_shm.
+Our second execution mode is the straight-line speculation mode or str, which means that we extend the sequential mode to also consider instructions up to a certain depth after a change in control flow (e.g., a branch or a jump) to expose their corresponding observations.
+We call this second contract str_shm.
+5.2. Hardware Side
+We now want to show that each of our hardware mechanisms satisfy one of these contracts, i.e., that they do not leak more information to an attacker that what is stated in the contract.
+In our case, we start from a design that satisfies Strong Microarchitectural Isolation.
+We also make sure to eliminate on-core microarchitectural state that could entangle state related to enclave private memory, shared memory and the attacker (see Section 6 where we bypass the L1-D cache).
+This means that the attacker is only able to observe the microarchitecture we explicitly share, i.e., the shared LLC partitions.
+As a result, the only architectural states the attacker is able to reconstruct through microarchitectural observations are the addresses of speculative shared memory accesses.
+Note that per our threat model described in Section 3 and following previous work, we artificially make the assumption that the attacker is not able to collect timestamps for the accesses it observes.
+For our first mechanism, SpecSafe (see Section 4.4), the program is executed following the out-of-order semantics of our processor, but we make sure to not touch any shared microarchitecture until a shared memory access has reached the head of the ROB and so is non-speculative.
+As a result, Citadel with SpecSafe verifies seq_shm.
+Our second mechanism is the burst mode (see Section 4.5).
+Here, the hardware was precisely designed to implement (see Section 4.5) straight-line speculation, and as a result obeys str_shm.
+Listing 1: Pathological Insecure Code for Burst Mode
+jal far_away_function
+lw a0, 0(secret_address)
+Listing 2: Assembly Snippet for the SHA256 library
+csrwi 0x802,BURST
+.loop1:
+[...] # Straight-line code with
+shared memory accesses
+bne a3,a6,.loop1
+csrwi 0x802,NO_BURST
+5.3. Software Side
+We now need to analyze if our programs can enforce SNI with regard to these contracts.
+In our case, SNI reduces to the fact that the program does not expose more secret data, when considered under a given security contract than when run sequentially, i.e., under seq_shm.
+This is straightforward for our first contract, as the contract is exactly seq_shm.
+Putting everything together, this means that any enclave program secure when run sequentially (i.e., that does not explicitly leak its secrets) is also secure when run in Citadel with SpecSafe.
+This is not true for all program for our second contract str_shm.
+Indeed, we can craft a pathological code snippet (see Listing 1) that would systematically leak a secret under str_shm while secure when run sequentially.
+Hence, burst mode can be harmful, and it is the responsibility of the programmer to make sure that they use burst mode in a safe context, by analyzing the code under the simple straight-line speculation model.
+Fortunately, we only use burst mode on small code snippets and are able to carry out the required program analysis by hand.
+This is possible as burst mode is enabled and disabled by writing to a special CSR with a csrwi instruction.
+This instruction cannot be performed speculatively and acts as a speculation barrier, hence it can be used to constrain the speculative flow control within clear bounds.
+Similar to previous work (guarnieri2021hardware, ; guarnieri2020spectector, ) we also make the assumption that when entering burst mode, the initial state of the microarchitecture touched by the code is independent of any secrets.
+This is reasonable as csrwi already flushes the processor pipeline, predictors are deactivated and the TLB and caches are easy to selectively warm up.
+The rest is left to the responsibility of the code before entering Burst mode.
+SHA256 Example We annotate a LOAD macro in a SHA256 library and show the corresponding assembly snippet in Listing 2 (full assembly in supplemental material).
+The performance impact of this simple annotation can be seen in Section 7.1.
+Here, the assembly shows us that the code is enclosed between two csrwi, which bound speculation.
+Furthermore, one branch is present, but is immediately fenced by a csrwi.
+It is fenced as csrwi is a system instruction that cannot enter the reorder buffer until the reorder buffer is empty.
+As a result, for that specific code snippet, the speculative control flow, and as a result, the exposed architectural states are the same for str_shm as for seq_shm, which is sufficient to say that the code snippet realizes SNI with regard to str_shm.
+memcpy Example We minimally annotate an implementation of memcpy to take advantage of burst mode when copying data from shared memory to enclave private memory (full C code in in supplemental material).
+More complex optimizations could use loop unrolling and annotate the second while loop.
+Performance results can be seen in Section 7.1.
+We show the details of the assembly for the code snippet we annotate in Figure 7.
+Here, the program analysis to verify SNI under str_shm is more complex as some shared memory accesses are performed speculatively.
+Nevertheless, by our definition of SNI, we only care about programs and inputs that lead to a secure execution (i.e., that does not leak secrets) under the seq_shm contract.
+That means that we can assume that registers a4 and a5 never hold any secret values under seq_shm as they are exposed to the attacker per our contract, through the shared memory accesses.
+Furthermore, at most one more run of .loop2 is executed speculatively.
+Hence, the only new values an attacker can learn are \$a4+8 and \$a5+8 with \$a4 and \$a5 the last values taken by the two registers under seq_shm and these values are not secret.
+Finally, which of these two values are exposed is public as we assume the initial microarchitectural state touched by the program to be public and no secrets are touched.
+Program Analysis Tools Finally, for more complex pieces of code, static analysis could be done manually or using many existing tools: e.g..
+oo7 (wang2019oo7, ), Kasper (johannesmeyer2022kasper, ), Kleespectre (wang2020kleespectre, ), Spectector (guarnieri2020spectector, ).
+Note that the usual critique and limitations of these tools are that 1) most code bases are too big to analyze without producing a lot of false positives 2) they make simplifying assumptions on the speculation model.
+Because Citadel also leverages hardware isolation, both of those limitations are not concerns.
+The code snippets we analyze are small and the observation and execution modes are simple, well-defined, and enforced by the hardware.
+Hence, these works are well-suited to our needs.
+We leave the automated analysis of programs by adapting these tools in the context of Citadel as future work (i.e., to support RISCV assembly).
+6. A few technical details
+We now elaborate on a few technical issues we faced while prototyping our machine.
+We also give a brief overview of the software infrastructure we have implemented in order to boot Linux and run enclaves from user programs.
+Speculative Transmitters in the L1 D Cache When enabling shared memory, the core-private part of the memory system, including the L1 D cache, is now shared between trusted memory accesses and untrusted or shared memory accesses.
 While shared memory cannot be accessed speculatively, trusted memory can.
@@
 As a result, a shared L1 D can be leveraged by an attacker to create a speculative transmitter and mount a Spectre-style attack.
-To our knowledge, we are the first to identify this subtle attack vector.
-Note that this doesn’t only apply to shared cache sets, but it also applies to contention over other shared resources in the L1 D cache (including the pipeline to access the SRAM).
-Partitioning the L1 D cache would be too costly, instead we bypass the L1 D cache for untrusted or shared accesses and directly send the requests to the LLC.
-4.5. Tagged TLB for Correct Address Translations
-When using the dual page table mechanism, the core-private part of the memory system is shared between trusted memory accesses and untrusted memory accesses.
-For functional correctness, we need to make sure that addresses are translated correctly for each domain.
-This small microarchitectural structure accelerates address translation by caching intermediate results used by the page walker.
-Unfortunately, only a few upper bits of the virtual address are used at the beginning of a page walk, easily resulting in aliasing across the two domains.
-We tag the Address Translation Cache by adding a bit to each entry indicating if the cached information applies to the trusted or untrusted memory address translations.
-Aliasing can also occur in the TLB if the trusted virtual memory range is smaller than the granularity of a virtual memory page (which can be if we use Mega or Giga Pages).
-We also tag the TLB entries to avoid address mistranslations.
-4.6. Miscellaneous
-We statically partition MSHRs between two cores and place a fair round-robin arbiter at the entry of the LLC.
-Because these two mechanisms only marginally affect the latency of the LLC for a 2-core design, we observe negligible ($\approx$ 0.1%) performance impact.
-We also expose controls for speculation and predictors to the software through Control Status Registers (CSRs).
-Specifically, a program is able to deactivate speculation all together (every instruction waits to reach the head of the ROB), delay the issuing of shared memory accesses to the head of the ROB, deactivate the training of predictors (branch predictors and return address stack) and/or their use (and then it falls back to the predictor $pc+4$).
-4.7. Building the Software Infrastructure
-We are interested in evaluating our platform on a Linux-ready system and for programs that one would be interested in putting in an enclave (e.g.
-cryptographic libraries).
-This means a lot of software needs to be developed to be able to launch enclaves from a Linux user mode process.
-First we implement the SM at the most privilege level of RISC-V using a mix of assembler and C and add the necessary pieces of a micro-kernel required to boot Linux on top of it.
-We also implement what we call a Mini-SM which is a downsized copy of the SM private to each enclave.
-This avoids contention over most SM resources between enclaves and untrusted programs.
-To protect the SM against transient execution attacks, we disable speculation during SM execution.
+Partitioning the L1-D cache would have a significant performance impact.
+Simply placing an arbiter in front of the L1-D arrays (which is very latency-sensitive) would have a much harsher overall impact on performance than for the LLC.
+At the same time, accesses to shared memory are typically not expected to be temporally reused (see Section 7.1), limiting the utility of the L1-D for shared memory anyway.
+For all these reasons, we decide to simply bypass the L1 D cache for enclave accesses to shared memory.
+Because the L1 is the only core-private microarchitectural state entangled with the adversary state through a coherence protocol, we do not need to worry about other core-private microarchitectural structures.
+Miscellaneous Security Specifics We statically partition MSHRs between the two cores, implement a fair round-robin arbiter at the entry of the LLC (not implemented in (bourgeat2019mi6, )), and arbiter the TLB requests (ignored and omitted in (bourgeat2019mi6, )).
+Because these mechanisms only marginally affect the latency of the LLC for a 2-core design, we observe negligible ($\approx$ 0.1%) performance impact in Citadel for insecure applications.
+Reconfigurable LLC Partitioning An obvious bottleneck in our design is the static cache partitioning inherited from MI6.
+Dividing our 1MB LLC in 64 equal-size 16KB partitions (see Section 4.1) is a terrible idea as these are smaller than the 64KB L1 D+I caches.
+Because our caches are inclusive, an enclave or a program that resides in one memory region cannot even fully use the L1 caches.
+MI6 circumvented that issue by assuming and simulating (but not fitting on FPGA) a 16-core design with a larger 16MB LLC.
+We build a mechanism to reconfigure the LLC partitions, improving enclave and other programs’ performance, regardless of the number of cores or the size of the LLC.
+This costly operation is only required before launching a new application or enclave with a different memory requirement.
+At the hardware level, we use a fine-grain set-partitioning scheme similar to (townley2022composable, ; saileshwar2021bespoke, ) to support our 64 partitions.
+Because for a given memory access, the memory region IDs are directly accessible through physical memory address, our scheme requires minimal changes in the LLC, no changes to the rest of the memory hierachy and supports shared memory between security domains.
+We also implement a fine-grain flushing mechanism by adding a simple piece of logic connected to the cache hierarchy at the same level as the DRAM, but which always sends back zeroes.
+This means that flushing a small LLC partition is much faster than flushing a big one.
+At the software level, the OS can request a reconfiguration through a SM call, which will ensure the LLC is not put in an incoherent state.
+The main challenge is to ensure that no core will emit requests to the modified regions during the operation.
+For the operation to proceed, the second core will have to perform a SM call to enter a loop located in a region unaffected by the reconfiguration.
+This means that an LLC reconfiguration is not possible if an enclave is running, some SM regions cannot be affected by the reconfiguration (checked by the SM).
+The SM will also check that the mapping is correct and fits within the LLC.
+Building the Software Infrastructure The overall software architecture and privilege hierarchy can be seen in Figure 1.
+We implement the SM at the highest privilege level of RISC-V using a mix of assembler and C.
+We also implement what we call a Mini-SM, which is a downsized copy of the global SM private to each enclave.
 Second, we develop a Linux Kernel module that allows user programs to request SM calls and set up enclaves, keeping Linux responsible for the resource management.
 Finally, we implement a Secure Bootloader and an end-to-end attestation mechanism that makes it possible to uniquely identify an enclave, establish a shared secret with it and exchange encrypted private data.
-Our attestation mechanism makes it possible for an enclave to be placed anywhere in physical memory and still have consistent identity through the measurement of the direct mapping between virtual addresses and memory content without measuring physical addresses.
-The unique mapping is attested and enforced by the SM during enclave setup.
-More implementation details can be found in the evaluation section (see section 6).
-5. Citadel 2.0: Improving Performance
-Our first baseline successfully achieves security.
-In this section, we show several mechanisms to improve its performance.
-First.
-fixed LLC partitioning á la MI6 destroys performance for our 2-core machine.
-We show a reconfigurable LLC partitioning scheme and a fine grain flushing mechanism.
-Furthermore, disabling speculative shared memory accesses makes software requirements especially light, but provides limited performances.
-We show a relaxed speculative model where we enable pipelined accesses to shared memory but disable predictors and hard-code them to $pc+4$.
-When this mode is enabled, the program needs to verify more requirements, i.e., that no secret is leaked when speculating $pc+4$.
-We show that this new mode improves performance for code segments with burst shared memory accesses, but it hampers performances in the general case.
-We observe that this is due to the TLB being trashed by misspeculations across functions context.
-This is an effect that was studied (in a nonsecurity context) in the early 2000s(onur1, ; onur2, ; onur3, ): misspeculations can induce large overhead not only due to the delay in executing the right path, large delays are induced due to the trashing of the TLBs.
-Finally, we show how insights from building the SM make it possible to securely optimize context switch and enable frequent interrupts.
-5.1. Evaluation-Guided Optimisations
-Our secure baseline runs on FPGA with a Linux environment (see section 6 for our full configuration) which makes it possible to analyses performance bottlenecks on real programs.
-For this section, we use an elliptic curves (Ed25519) cryptographic library as a running example (see section 6 for more details).
-At a high level, our enclave interactively receives messages to sign from a benchtest program using shared memory.
-Out-of-the box, we see an underwhelming overhead of 120% overhead compared to directly (and insecurely) linking the cryptographic library to our benchtest program.
-5.2. Reconfigurable LLC Partitioning
-An obvious bottleneck in our design is cache partitioning.
-Sanctum and MI6’s mechanism (i.e.
-dividing the LLC in 64 equal size slices to match the 64 DRAM regions) is a terrible idea for our system.
-Our 2-core prototype has a 1MB LLC (see 5), which divided by 64 gives us 16KB LLC slices.
-This is smaller than the 64KB L1 D+I caches.
-Because our caches are inclusive, this means an enclave or a program that resides in one memory region cannot even fully use the L1 caches.
-MI6 circumvented that issue by assuming a 16-core design with a large 16MB LLC.
-We build a mechanism for the OS to reconfigure the LLC partitions to improve enclave and other program’s performance.
-This costly operation is only required when changing the LLC layout, which only happens before launching an application or an enclave with an important memory footprint.
-The OS requests a reconfiguration through an SM call while no enclaves are running.
-The SM will first check that the new partitioning is valid, doesn’t affect regions assigned to the SM.
-It then waits for other cores to flush their L1 caches and to enter a blocked state in the SM, making sure no requests will be sent to the modified regions during reconfiguration (which could put the cache in an incoherent state).
-The SM will then interact with the hardware to flush the LLC slices that are modified and set the LLC in its new configuration.
-At the hardware level, we now design a practical reconfigurable LLC partitioning scheme.
-We cannot use way partitioning (kiriansky2018dawg, ) as the LLC has only 16 ways, and we need up to 64 slices, and Ironhide’s (ironhide, ) strategy assumes limited reconfiguration.
-We propose a new, simple, configurable set partitioning scheme that only requires minimal changes in the hardware.
-The configuration for each LLC slice (base and bound of the set-index range) is stored in a big register (1216 bits) called the set-index range table.
-The SM can modify the entries of the set-index range table by sending MMIO requests to the cache, which is seen as a new device.
-When the LLC receives a request, it will compute the new set index for the corresponding address.
-We represent the transformation used in Figure 4.
-First, we extract a region ID from the address upper bits and obtain the corresponding index range base and size by looking up the set-index range table 1.
-We then extract the address original set index, located in the middle of the address, and tweak it.
-We compute the original set-index value modulo the size of the index range 2 and add the base index 3.
-This effectively restricts the new set index to the corresponding index range, partitioning the LLC 4.
-Note that we must extend the length of the tags used by the LLC to include the entire original index in case an index range is of size 1 and all addresses from the same region map to the same set.
-We observe on our cryptographic library that we reduce overheads from 120% to 19.4% by using a non-uniform partition for the LLC. More details can be found in Section 6.5.
-Figure 4. LLC set indexing scheme. The function takes an address as an input and output a set index in the LLC slice allocated to its memory region.
-LLC fine-grain flush To go hand-in-hand with our reconfigurable LLC partitioning, we also want to be able to selectively flush parts of the LLC.
-We create a Zero Device in the upper physical address space.
-It is implemented as a simple piece of logic connected to the cache hierarchy at the same level as the DRAM, but that always sends back zeros.
-The trick is that the address space covered by the zero device has the same size as the DRAM and is located at addresses that will collide with DRAM addresses in the cache.
-Once the SM has placed the machine in a correct waiting state (where all the cores have flushed their L1 and are waiting in the SM), it only needs to access an eviction set in the zero-device address space to flush the cache slice.
-This means that flushing a small LLC slice is much faster than flushing a big one.
-We look at the performance of our flushing mechanism in the evaluation section (see 6.5).
-5.3. Pipelined Shared Accesses
-At this step, we observe that our test benchmark still present a 19.4% overhead compared to our insecure benchmark.
-We decide to measure the performance of different sections of the code to find and analyze the bottleneck.
-We identify the function creating the slowdown to be located inside the SHA256 library.
-This function performs burst accesses to shared memory to load the message to be signed.
-This behavior is exacerbated by the compiler, trying to optimize the code by placing all the load instructions to shared memory one after the other.
-Indeed, with our current mechanism, memory accesses to shared regions, first cannot use the L1 and have a 10 cycle latency, second cannot be pipelined as each instruction needs to wait to reach the head of the ROB to be sent.
-This is especially costly during burst accesses.
-The LLC supporting 16 outstanding requests and having a 10 cycles latency, 16 memory accesses should be pipelined to take a total of $10+16=26$ cycles.
-Without pipelining, they will now take $10*16=160$ cycles.
-As a result, we designed a mode where we temporarily change the speculation model to allow shared memory burst accesses (as pipelining is a form of speculation).
-Nevertheless, we don’t want to do so while enabling any type of speculation.
-Indeed, we don’t want our program to speculatively jump to a spectre-like gadget and transmit secrets to our attacker.
-To avoid this, we tweak the speculation model to allow pipelined accesses to shared memory but disable the use and training of any predictors that could result in the speculative execution of rogue code.
-Specifically, we disable the Branch Target Buffer (BTB), the Branch History Table (BHT) and the Return Address Stack (RAS).
-Effectively, that guarantees that the hardware will always predict that the next program counter to be “$pc+4$” and avoid speculative execution of arbitrary code.
-We call this straight-line speculation.
-Modifying the speculative model also modifies the hardware software contract.
-The programmer now needs to ensure that when this mode is enabled, no speculative transmitter can be reached by the program when doing straight-line speculation.
-Static analysis can be done manually or by reusing many existing tools: e.g.
-oo7(wang2019oo7, ), Kasper(johannesmeyer2022kasper, ), Kleespectre(wang2020kleespectre, ), Spectector(guarnieri2020spectector, ).
-Note that the usual critique and limitation of these tools are that 1) the code bases are too big to analyze without producing a lot of false positives 2) they make simplifying assumptions on the speculation model.
-Interestingly, both of those limitations are not concerns in the context of Citadel’s burst mode because the code snippets we analyze are small (e.g.
-boiling down to a memcpy function for instance) and the speculation model is well-defined, simple and enforced (only speculates PC+4).
-Hence, these works are well-suited to our needs.
-We consider adapting these tools to our more simple setup as future work.
-We add two new configurations in our software-exposed configuration registers to enable and disable this “burst” mode.
-The right way to use this mode is, like in our case, during burst accesses to shared memory (e.g.
-a memcpy function moving data from public memory to private enclave memory).
-The programmer only needs to identify these series of burst access and guard them with instructions activating and deactivating burst mode.
-When doing so on our cryptographic benchmark, we observe that overhead falls from $19.4\%$ to $2.7\%$.
-More results are presented in the evaluation section.
-5.4. Relaxed Flushing Policies
-Because we build the entire software infrastructure required to run enclaves, we were able to observe that some security domains, especially the enclave and some parts of the mini-SM, do not have secrets to keep from each other.
-Specifically, the mini-SM never manipulates sensitive data.
-As a result, the mini-SM does not need to flush microarchitectural state when context switching to enclaves.
-This reduces the cost of mini-SM call and opens the door to the use of frequent software or timer interrupts.
-To disable flushing selectively, we modify the hardware to deactivate systematic flushing and exposed to the SM a register to trigger the flush.
-We write the SM to trigger the flushing mechanism on required context switches, but not on most traps and SM-calls from the enclave.
-We evaluate the performance implication of our mechanism in Section 6.6.
 7. Evaluation
-6.1. Experimental Setup
-Frontend
-2-way superscalar, 256 entries BTB, 8 entries RAS, tournament predictor
-Reorder Buffer
-64 entries, 2-way insert/commit
-Reservation Stations
-2 stations for ALU (16 entries), 1 for FPU (16 entry), 1 for memory (16 entries)
-L1 (I/D)
-32KB, 8 ways, 8 outstanding requests
-Load-Store Unit
-24 LdQ entries, 14 StQ entries, 4 SB entries
-LLC (L2)
-1MB, 10 cycles latency, 16 ways, max 16 outstanding requests
-Memory
-120 cycles latency (24 outstanding requests)
+Experimental Setup
+Frontend 2-way superscalar, 256 entries BTB, 8 entries RAS, tournament predictor
+Reorder Buffer 64 entries, 2-way insert/commit
+Reservation Stations 2 stations for ALU (16 entries), 1 for FPU (16 entry), 1 for memory (16 entries)
+L1 (I/D) 32KB, 8 ways, 8 outstanding requests
+Load-Store Unit 24 LdQ entries, 14 StQ entries, 4 SB entries
+LLC (L2) 1MB, 10 cycles latency, 16 ways, max 16 outstanding requests
+Memory 120 cycles latency (24 outstanding requests)
 Figure 8. Citadel’s microarchitectural configuration.
 Citadel’s processor is a modified out-of-order 2-core based on Riscy-OOO (zhang2018composable, ) and MI6 (bourgeat2019mi6, ).
-The parameters used for the cores and the memory subsystem are given in Figure 5.
+The parameters used for the cores and the memory subsystem are given in Figure 8, and kept small to keep the synthesis and simulation time small.
 Citadel runs on an AWS EC2 F1 instance, equipped with a Virtex UltraScale+ FPGA.
@@
 We run Linux Kernel 6.2.0.
-Unless mentioned differently, all performance numbers are collected by running the benchmark on the FPGA using performance counters and an average of ten runs.
-6.2. Implementation Data and TCB Breakdown
-Component
-LOC Size
-Processor (bourgeat2019mi6, ) (+ modification)
-$\sim 60$ KLOC (+ 250 ) N/A
-Bootloader
-5236 55KB
-Security Monitor
-9287 97KB
-Mini-SM
-2301 5.6KB
-Linux
-$\sim 20$ MLOC 157MB
-SM Kernel Mod.
-507 505KB
-TCB Software
-14523 152KB
+All performance numbers are collected by running the benchmarks on the FPGA using performance counters and an average of ten runs.
+Implementation Data and TCB Breakdown
+Component LOC Size
+Processor (bourgeat2019mi6, ) (+ modification) $\sim 60$ KLOC (+ 250 ) N/A
+Bootloader 5236 55KB
+Security Monitor 9287 97KB
+Mini-SM 2301 5.6KB
+Linux $\sim 20$ MLOC 157MB
+SM Kernel Mod. 507 505KB
+TCB Software 14523 152KB
 Table 1. Software Component Size and TCB Breakdown. Mini-SM is a subset of the SM code. Lines in red correspond to elements excluded from the TCB.
@@
 Final TCB size needs to include the enclave’s binary, which depends on the application.
+7.1. Evaluating Secure Shared Memory
+Microbenchmark on memcpy We evaluate two different setups, one with the function as is (only using SpecSafe) and the other one when minimally annotated for Burst mode (full C code in supplemental material and assembly snippet in Figure 7).
+We normalize the results to using memcpy to and from private memory (not using our mechanisms) in Table 7.1.
+We observe that out-of-the-box, SpecSafe shows dramatic overhead, especially when the data is not cached in the LLC with a 2231% overhead.
+On the other end, we show that Burst mode is quite efficient and reduces the overhead to 245% and 62% depending on whether the data is cached in the LLC.
+We also try different configurations where we start with a cold LLC (all loads will go to main memory), start with a warm LLC (the OS has accessed it previously or the data just arrived from the network for instance) or an insecure setup, where we have enabled the L1 for shared memory accesses.
+Our results confirm that access latency has a huge impact on the performance of SpecSafe, which is mitigated by Burst Mode.
+Not shown in the table, we observe that when moving data from private memory to shared memory, the overhead is 48% in all cases (both mechanisms and with cold/warm LLC).
+Indeed, store operations to shared memory are already pipelined with SpecSafe and Burst mode does not improve the performance.
+We also measure the overhead of straight-line speculation to be around 33% by enabling Burst mode while using memcpy to and from private memory.
+Finally, looking at performance counters, we observe that the  20% overhead left can be attributed to address translations, which are already cached when performing private memory copy.
+L1-D Bypass These results also highlight the reasonable performance impact of bypassing the L1-D for shared memory accesses, especially when using Burst mode with a warmed up LLC.
+Note that we can also make a qualitative argument for bypassing the L1-D as shared memory accesses usually do not involve re-access (computing on shared memory could be insecure as the access pattern is leaked) and getting the output data to the LLC as fast as possible makes a lot of sense as we want it to be accessed by another security domain.
+SpecSafe Burst Mode
+No LLC warm-up 23.317 3.458
+LLC warm-up 3.665 1.621
+LLC warm-up + L1 2.466 1.556
+Table 2.
+memcpy micro-benchmark.
+Copy of 1500 bytes from shared memory to private memory.
+Normalized over memcpy from an to private memory with speculation enabled.
+We experiment with warming up the LLC and enabling the L1 for shared memory accesses (insecure).
+{NiceTabular}
+[…]
+[…]
+Citadel
+(SpecSafe + Burst) 1.027 1.464
+[…]
+SpecSafe Only 1.194 4.973
+[…]
+Burst Only 1.654 4.959
+Table 3.
+Normalized runtime for signing and hashing three thousand random messages sent by the OS to an enclave.
+Runtime is normalized to the insecure baseline where the library is directly linked to the testbench program.
+We compare Citadel with two configurations where we are not using burst mode or insecurely using burst mode everywhere.
+Cryptographic Library We now evaluate our mechanisms for secure shared memory on a realistic application by wrapping a cryptographic library in an enclave (see details in Section 7.2).
+We measure the time it takes for the enclave to perform $3000$ cryptographic operations (signature or cryptographic hash) over random inputs provided by the OS.
+We compare three setups in Table 7.1.
+First, our final design that combines spec-safe shared memory + burst mode enabled around performance sensitive shared memory burst sequences (Citadel).
+Then Citadel without annotating code for shared memory burst sequences (Spec-Safe Shared Memory only).
+Finally, another (insecure) configuration where we annotate all the code to be run in burst mode (Burst only).
+We report the runtime normalized to the insecure baseline (directly linked library and no LLC partitioning).
+We observe an overhead of $2.7\%$ on the signature benchmark and $46.4\%$ for the hashing benchmark.
+This is because the hashing application accesses shared memory much more frequently compared to the signing application that is more computationally bounded.
+This difference between the benchmarks is also highlighted by our other results, where not annotating the code for burst shared accesses has a $397.3\%$ overhead for the hash application and a $19.4\%$ overhead for the signing application.
+These results show that the burst mode improves performance for code segments with burst shared memory accesses.
+Burst Mode in the General Case On our benchmarks, we observe that the performance overhead goes up to $65.4\%$ for the signing benchmark and $395.9\%$ for the hashing benchmark when (insecurely) activating burst mode for the whole enclave (Table 7.1).
+This overhead is due to the TLB being trashed by straight-line misspeculations across function contexts.
+Instead of speculating a jump to the return address using the Return Address Stack (RAS) at the end of the function, the processor will speculatively run the prologue of the next function in memory.
+We observe that many of these functions will try to deference pointers passed as arguments, triggering translations for arbitrary values in this erroneous context.
+This will evict useful cached translations from the TLB causing an excessive slowdown later on.
+This effect was already documented (in a nonsecurity context) in the early 2000s (onur1, ; onur2, ; onur3, ).
+We confirm this hypothesis by instrumenting the binary and observing that when the RAS is reactivated (but not other predictors), the number of TLB misses, page walks and the overall execution time goes down significantly.
+Note that we do not address this performance issue, but instead use burst mode for performance-critical code snippets, which also makes security analysis tractable (see Section 5).
 7.2. Evaluating the End-to-End Platform
-MicroPython Runtime We do a minimal port of MicroPython (microphython, ) inside one of our enclaves.
-We start with the bare-metal ARM implementation.
-We modify the configuration files and the linker script so our binary can run in the virtual memory layout provided by our enclave.
-We only need to implement a few functions by hand.
-For instance, we redirect the console to print to shared memory and need to write a bit of assembly for the setjmp and longjmp functions used for exception handling.
-We also add a cryptographic library, so the enclave can sign the console output using its own private keys recovered through the remote attestation mechanism.
-A hash of the standard output is extended on each character print and signed on the enclave exit.
-We cross-compile MicroPython using the RISC-V tool-chain, build without standard library, and generate position independent code.
-In total, we need to manually modify 128 LOC for the port and add the 5K LOC cryptographic library.
-The final compiled binary measures 180KB.
-With the enclave’s code added, the TCB is now 65,174 LOC, most of it (70%) being MicroPython.
-Figure 6. MicroPython Enclave Execution Breakdown.
-We use a simple primality test as a benchmark where a number is found to be prime if none of the numbers smaller than it can divide it.
-We test the primality of an arbitrary small prime number: $19,531$.
-The tests take an average of $4.1\times 10^{7}$ cycles to execute, with 51% of the time spent within the Python program.
-This means the enclave and attestation mechanism add 94% overhead for this small computation.
-We show a visual breakdown of the running time in Figure 6.
-We note that most of this overhead comes from hashing enclave code (25% of the total running time) but also generating the enclave keys and signing the certificate (10% of the running time).
-These two numbers are especially high since this code is run non-speculatively inside the SM.
-On another hand, hashing and signing the output inside of enclave takes a negligible amount of time (less than 1% of the total running time).
-Because this cost is fixed, it could definitely be amortized for bigger applications.
 Securing a Cryptographic Library We write a wrapper around an ED25519 RISC-V library (ed25519github, ).
-The wrapper makes it possible to keep the keys used by the library inside of the enclave and for the library functions to be accessed in a Remote Procedure Call style (RPC).
 We add a queue to shared memory, so untrusted applications can send requests to the library to generate keys (that will stay in enclave memory), or to sign messages using previously generated keys.
@@
 Nevertheless, context-switching in and out of an enclave is expensive as it requires flushing the microarchitectural state, which makes the slowdown more important in the re-entry scenario.
-These two secure solutions highlight an interesting tradeoff between resource occupation (one core and a DRAM region vs.
-only a DRAM region) and performance.
-Private Inference We evaluate the following scenario where a remote user wants to do perform some private inference task on a neural network evaluated on an untrusted server.
-We pace a simple neural network model inside of one of our enclaves.
-We use onnx2c (onnx2cgithub, ) to convert a ONNX file describing a handwritten digit recognition model (trained on MNIST) to standard C code and run private inference inside of our enclave.
+These two secure solutions highlight an interesting tradeoff between resource occupation (one core and/or a DRAM region) and performance.
+Private Inference We evaluate the following scenario where a remote user wants to perform a private inference task on a neural network evaluated on an untrusted server.
+We place a simple neural network model inside of one of our enclaves.
+We use onnx2c (onnx2cgithub, ) to convert a ONNX file describing a handwritten digit recognition model (trained on MNIST) to standard C code and run private inference inside our enclave.
 The user uses our remote attestation mechanism to verify the enclave identity and establish a shared secret with it.
@@
 It then sends back an encrypted output to the remote user.
-In total, we need to implement 227 lines of straightforward C, which is to be compared to the 1536 LOC of the neural network and the 410 LOC of the AES library and the 4386 LOC of ED25519 we now need to include inside of the enclave for encryption and key exchange.
+In total, we need to implement 227 lines of straightforward C, which is to be compared to the 1536 LOC of the neural network and the 410 LOC of the AES library and the 4386 LOC of ED25519 we also include for encryption and key exchange.
 With the enclave code added, the TCB is now 21K LOC and the final enclave binary measures 54KB.
-We now compare the performance between our setup and an insecure baseline where the model is directly evaluated outside of an enclave and where the LLC is not partitioned.
-We do consider that incoming traffic from the remote user is encrypted and needs to be decrypted first.
-In both setups, runtime is split 55% and 45% between AES operations and inference.
-We see negligible (0.76%) overhead for running a model inside of the enclave, with all of it coming from the necessity to copy the ciphertext in enclave private memory before performing decryption.
+We now compare the performance between our setup and an insecure baseline where the same benchmark (including decryption / encryption) is directly evaluated outside an enclave and where the LLC is not partitioned.
+In both setups, runtime is split 55% and 45% between AES and inference.
+We see negligible (0.76%) overhead for running a model inside the enclave, with all of it coming from the necessity to copy the ciphertext in enclave private memory before performing decryption.
 This is not surprising as the rest of the computation happens in enclave private memory and should not suffer from any slowdown.
-This doesn’t account for the time required to set up the enclave, which is a one-time cost but represents 23.3% of the cost of performing 1024 inference.
-6.4. Evaluating Secure Shared Memory
-We first demonstrate that naive public memory sharing is vulnerable to transient execution attacks by mounting an attack on an insecure version of Citadel where speculation on shared memory is allowed.
-The attack uses the LLC as a side channel and use a specter gadget located in enclave code to leak arbitrary enclave memory to untrusted code ruining on another core.
-Anecdotally, we were not able to successfully mount the attack on our secure prototype.
-Table 2.
-Normalized runtime for signing and hashing three thousand random messages sent by the OS to an enclave.
-Runtime is normalized to the insecure baseline where the library is directly linked to the testbench program.
-We compare Citadel with two other configurations where the library code is not annotated for burst accesses to shared memory (- Burst) or all annotated (++ Burst).
-[…]
-[…]
-[…]
-[…]
-We then evaluate the perfromance overhead of our comunication mechanism.
-We use a similar setup as Section 6 but only measure time spent computing the cryptographic function.
-The enclave performs $3000$ cryptographic operations (signature or cryptographic hash) over random inputs provided by the OS.
-We compare three setups.
-First, our final design with burst mode enabled around shared memory burst sequences (Citadel).
-Then Citadel without annotating code for shared memory burst sequences (- Burst).
-Finally, another (insecure) configuration where we annotate all the code to be run in burst mode (++ Burst).
-Our results are presented in Table 6.4.
-We report the runtime normalized to the insecure baseline (directly linked library and no LLC partitioning).
-We observe an overhead of 2.7% on the signature benchmark and 46.6% for the hashing benchmark.
-This is because the hashing application accesses shared memory much more frequently compared to the signing application that is more computationally bounded.
-The hashing function is more affected by the loss of the L1 cache.
-This difference between the benchmarks is also highlighted by our other results, where not annotating the code for burst shared accesses has an $397.3\%$ overhead for the hash application and a $19.4\%$ for the signing applicaiton.
-These results also demonstrate the efficiency of our burst mechanism when used properly.
-We also investigate the performance of using this mode on the whole enclave code instead of the limited code segment when the enclave communicates.
-On our benchmarks, we observe that the performance overhead goes up to $65.4\%$ for the signing benchmark and $395.9\%$ for the hashing benchmark.
-We investigate the origin of this overhead and realize that miss speculation due to the deactivation of the RAS generates a lot of random load that pollutes the TLB, triggers many page walks and terribly slows down the virtual memory subsystem.
-This is due to many small functions being linked one after the other in the binary.
-Some of these functions start with loads using generic registers as addresses (for instance when accessing a pointer passed as an argument to the function).
-When encountering a ret instruction, the processor will first speculate that the following instructions should be executed.
-In this speculative window, it will start running the instructions of the function linked right after and perform loads or stores to random addresses from registers containing garbage.
-We confirm this hypothesis by instrumenting the binary and observing that when the RAS is reactivated (but not other predictors), the number of TLB misses, page walks and the overall execution time go down significantly.
-As a result, we recommend a hybrid approach where the burst mode is only used when the enclave wants to communicate.
-Table 2.
-Normalized runtime for signing and hashing three thousand random messages sent by the OS to an enclave.
-Runtime is normalized to the insecure baseline where the library is directly linked to the testbench program.
-We compare Citadel with two other configurations where the library code is not annotated for burst accesses to shared memory (- Burst) or all annotated (++ Burst).
+This doesn’t account for the time required to set up the enclave, which is a one-time cost but represents 23.3% of the cost of performing 1024 inferences.
+MicroPython Runtime We do a minimal port of MicroPython (microphython, ) inside an enclave.
+The goal is to provide an attested Python runtime, with a signed transcript of the interactions.
+It serves as a small benchmark to demonstrate the expressivity and flexibility of the platform.
+We start from the ARM port and only need to modify a few of the MicroPython runtime functions.
+For instance, we redirect the console to print to shared memory and need to write a bit of assembly for the setjmp and longjmp functions used for exception handling.
+Moreover, a hash of the standard output is extended on each character print and the hash is signed on the enclave exit using a cryptographic library and the enclave’s own private keys (obtained through the attestation mechanism).
+In total, we modify 128 LOC of MicroPython for the port to our platform, and add the 5K LOC cryptographic library.
+With the enclave’s code added, the TCB is now 65,174 LOC, most of it (70%) being MicroPython.
+We test our attested runtime on a simple Python program: a simple primality test.
+We successfully run the test, and verify the signature of the transcript.
+7.3. Evaluating LLC Dynamic Partitioning
+We evaluate the advantage of dynamically partitioning the LLC on our cryptographic enclave benchmark.
+We compare three different LLC configurations: an insecure baseline where the LLC is not partitioned, a secure configuration where the LLC is statically partitioned into uniform partitions of 16KB (MI6 scheme) and a last configuration where we use our scheme to partition the LLC non-uniformly.
+For the non-uniform configuration, we split the 1MB of LLC as follows: we give 256KB to the enclave, 16KB to the SM, 128KB to the OS, 128KB to the region used for shared memory, 128KB to the region storing page tables, and 1KB each to every other region.
+We observe a 67.8% overhead for the uniform partitioning and 0.08% overhead for the non-uniform partitioning, highlighting the benefits of dynamic partitioning.
+This is especially important as untrusted applications also benefit from it.
+We also measure the time it takes to modify LLC partitioning.
+We measure that the operation takes on average 1M cycles.
+This is a very costly operation, but it only happens very infrequently.
+Among these, 98% of the cycles are spent flushing the LLC.
+As a result, changing the layout of fewer regions is much faster: we measure an average of 2.5K cycles to flush a single LLC set containing 16 ways.
+The bigger the number of LLC set indexes affected, the longer the operation will take.
+7.4. Takeaways for Commercial Systems
+Our performance evaluation gives us insights on the potential impact of hardening commercial systems against side channels and transient execution attacks.
+We demonstrated that hardware with both bare-metal strong isolation primitives and a controlled speculation mode, akin to our Burst mode, enables system developers to create libraries and abstractions to interact with shared memory with good performance, without taking the usual security risks related to speculated executions.
+These mechanisms might be even more relevant when considering commercial-grade cache-hierarchy with higher latencies.
+We envision that the same kind of controlled speculation mode is implementable in commercial machines, to give most of the benefits of speculation for easy-to-verify mostly-straight-line code when enclaves access shared memory, with a simple and user-understandable speculation model.
+Note that enclave memory accesses can have unrestricted speculation.
+While implementing such controlled speculation mode and hardware primitives on modern industrial processors would be more involved than for our prototype, we emphasize the small relative complexity of the features compared to the rest of the design (250 LOC out of 60K LOC).
+We also showed that the typical cost of LLC partitioning for isolation can be greatly mitigated with a simple design that allows dynamic resizing of the partitions.
+We believe that all performance optimizations presented will also improve commercial systems, albeit with different quantitative benefits.
+8. Related Work
+Mechanisms for Secure Speculation Many academic works (ainsworth2020muontrap, ; cheang2019formal, ; canella2019systematic, ; barber2019specshield, ; fustos2019spectreguard, ; yan2018invisispec, ; yu2019speculative, ; saileshwar2019cleanupspec, ; weisse2019nda, ; khasawneh2019safespec, ; loughlin2021dolma, ; ainsworth2021ghostminion, ) present hardware mechanisms or hardware-software co-design to defend against these attacks.
+They sometimes aim for a stronger security property than us (speculative constant time) but do so within threat models that restrict adversarial power to control or observe only a subset of microarchitectural state (e.g., only consider cache side channels) and only consider specific forms of speculation (e.g., branch predictors cannot be updated speculatively).
+These threat models could be seen as indirectly building assumptions on the microarchitectural isolation and the speculation model, making these mechanisms potential add-ons to a Citadel that satisfies these isolation requirements.
+As a result, we see our work as complementary.
+We note that very few of these mechanisms have been implemented in hardware, i.e., FPGAs or silicon.
+Many attacks that break these defense mechanisms rely on exploiting details that are often glossed over in simulation (like not faithfully simulating the memory hierarchy, for instance).
+Enclave Platforms Sanctum (costan2016sanctum, ) introduces the concept of a privileged security monitor by decoupling enclave management functionalities from hardware mechanisms.
+Our platform draws many concepts from Sanctum such as the dual page table and the memory isolation mechanism.
+Sanctum also considers some cache side channels in its threat model.
+Nonetheless, Sanctum is not concerned with speculative side channels and no functioning enclave system was ever built.
+Other platforms such as Keystone (lee2020keystone, ), CURE (bahmani2021cure, ), or (enclavecaches, ) address cache side channels, but ignore other side channels and only consider in-order processors.
+Many other academic platforms exist (ferraiuolo2017komodo, ; feng2021scalable, ; weiser2019timber, ; mccune2008flicker, ; azab2011sice, ; xia2013architecture, ; deng2014equalvisor, ; brasser2019sanctuary, ; sun2015trustice, ; koeberl2014trustlite, ; aegis, ; xom, ), but none of them consider speculative side channels or transient execution attacks to the extent of Citadel.
+MI6 Citadel builds on the open-source code provided in MI6 (bourgeat2019mi6, ) and reuses many of its hardware mechanisms.
+The MI6 paper focused on evaluating the performance of these mechanisms on an FPGA by running SPEC benchmarks directly on top of (insecure) Linux and using Linux system calls as hooks to project the cost of a potential context switch to the SM.
+The platform (likely missing the necessary software) did not evaluate a full enclave scenario (neither did it guarantee any security when running SPEC).
+In this paper, on top of implementing secure shared-memory and several mechanisms to improve performance and configurability, we have built and evaluated an end-to-end hardware-software platform, focusing on enclaves that run security-aware workloads and uncovering practical insights for enclave platforms.
+Note that Citadel’s performance numbers for SPEC would be similar to MI6’s evaluation as Citadel’s LLC can be configured to match the partition size used in (bourgeat2019mi6, ).
+Enclave Execution Models Recent works have tried to make memory sharing and enclave execution more extensible and configurable (van2023cheri, ; li2021confidential, ; lee2022cerberus, ; park2020nested, ; steinegger2021servas, ).
+Those works are focusing on architectural capabilities and isolation, but ignore security challenges introduced by microarchitectural side-channels and transient execution attacks.
+Under such a threat model, even basic communication primitives like Citadel public shared memory with the OS are not straightforward to implement efficiently and securely.
+9. Conclusion
+This paper advocates for processors to incorporate microarchitectural isolation primitives and mechanisms for controlled speculation, enabling different execution modes.
+These modes can restrict what is exposed to an attacker, effectively balancing performance and program-analysis complexity.
+We introduce two mechanisms to securely share memory between an enclave and an untrusted OS in an out-of-order processor.
+We show through security analysis and performance evaluation, that these two modes are complementary, achieving speculative non-interference with a reasonable performance impact, while requiring minimal code annotation and simple program analysis doable by hand.
+Our prototype, Citadel, is a multicore processor running on an FPGA, booting untrusted Linux, and supporting comprehensive enclave capabilities, such as shared memory, and remote attestation.
+To our knowledge, Citadel is the first end-to-end enclave platform to run secure applications, such as cryptographic libraries or small private inference workloads, on a speculative out-of-order multicore processor.